### v3

# A Fully Data-Driven Value Iteration for Stochastic LQR: Convergence, Robustness and Stability

Abstract.
Unlike traditional model-based reinforcement learning approaches that estimate system parameters from data, non-model-based data-driven control learns the optimal policy directly from input-state data without any intermediate model identification. Although this direct reinforcement learning approach offers increased adaptability and resilience to model misspecification, its reliance on raw data leaves it vulnerable to system noise and disturbances that may undermine convergence, robustness, and stability. In this article, we establish the convergence, robustness, and stability of value iteration (VI) for data-driven control of stochastic linear quadratic (LQ) systems in discrete-time with entirely unknown dynamics and cost. Our contributions are three-fold. First, we prove that VI is globally exponentially stable for any positive semidefinite initial value matrix in noise-free settings, thereby significantly relaxing restrictive assumptions on initial value functions in existing literature. Second, we extend our analysis to settings with external disturbances, proving that VI maintains small-disturbance input-to-state stability (ISS) and converges within a small neighborhood of the optimal solution when disturbances are sufficiently small. Third, we propose a new non-model-based robust adaptive dynamic programming (ADP) algorithm for adaptive optimal controller design, which, unlike existing procedures, requires no prior knowledge of an initial admissible control policy. Numerical experiments on a “data center cooling” problem demonstrate the convergence and stability of the algorithm compared to established methods, highlighting its robustness and adaptability for data-driven control in noisy environments. Finally, we apply the method to dynamic portfolio allocation, demonstrating its practical relevance outside traditional control tasks.
aDepartment of Mechanical Engineering, University of New Mexico, Albuquerque, New Mexico 87110 USA (e-mail: lcui@unm.edu).
bDepartment of Electrical and Computer Engineering, Tandon School of Engineering, New York University, Brooklyn, NY 11201 USA (e-mail: zjiang@nyu.edu).
cDepartment of Mathematics, Courant Institute of Mathematical Sciences, New York University, Manhattan, NY 10012 USA (e-mail: petter.kolm@nyu.edu; g.macqueron@nyu.edu).
Corresponding author (g.macqueron@nyu.edu).
Keywords: Adaptive dynamic programming, data-driven control, reinforcement learning, stochastic discrete-time linear quadratic regulator, small-disturbance input-to-state stability, value iteration.
©2026 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
1. Introduction
Machine learning (ML) and learning-based methods have become integral to science, engineering, and technology, significantly impacting control theory while introducing new challenges and open problems in the field. Unlike conventional model-based control approaches, which rely on mathematical models derived from first principles, learning-based control learns directly from data and optimizes control strategies. This enables learning-based control to handle complex systems with flexibility and adaptability beyond what manual tuning and traditional model-based strategies offer, thereby enhancing its practical applicability and value. Researchers commonly employ reinforcement learning (RL) techniques to develop learning-based control methods, allowing agents to interact with their environments, make decisions, and minimize cumulative costs [56].
Recent research examines RL algorithms within settings characterized by continuous states and actions, employing the linear quadratic regulator (LQR) as a standard benchmark to evaluate the convergence and robustness of these algorithms [2, 18, 1, 22]. Since Kalman introduced the LQR in 1960 [33], subsequent work has demonstrated its theoretical tractability and broad applicability in the controller design of real-world engineering systems [55, 25, 16].
Commonly, RL methods for LQR build upon policy iteration (PI) and value iteration (VI). PI-based approaches require a stabilizing control policy at the beginning of the learning process [5, 38, 42]. However, this requirement is restrictive in practice. When the system model is not fully accessible, identifying an initial stabilizing controller becomes a significant challenge. Additionally, each PI iteration requires solving a matrix equation, making the method computationally expensive in higher dimensions [24, 37, 14, 26, 8]. In contrast, VI does not require an initial stabilizing control policy or the solution of a matrix equation at each step. Instead, VI starts from an arbitrary positive definite matrix that serves as an initial approximation of the value function [6, 10, 12]. Within a noise-free environment, and under the assumption of detectability and stabilizability, VI is guaranteed to converge to a stabilizing and optimal controller, provided that each iteration of VI is computed exactly [7, 10]. These advantages make VI one of the most widely used and thoroughly understood algorithms for solving discounted Markov decision processes (MDPs) [49]. Consequently, VI finds broad application in various engineering domains, including wheel-legged robots [17], sensory-motor control [16] and autonomous vehicles [25].
Beyond the idealized case of noise-free environments, a practical learning-based control algorithm must also be robust to perturbations arising from noisy measurements, real-world data, and numerical errors. In data-driven control scenarios, where exact system dynamics are unknown, approximating the value function using sampled data introduces errors due to noise or insufficient data. Consequently, it is essential to determine whether the VI method remains robust to such errors in the learning process. In particular, when subjected to noise and errors, whether VI converges to a near-optimal solution? While VI has been extensively studied, its convergence and robustness under these conditions remain unresolved.
Recent work on exact VI establishes exponential convergence only under restrictive assumptions on the initialization. In particular, [43] proves exponential convergence when the initial value lies in a semidefinite cone around the optimum, and [52] derives local exponential convergence in a neighborhood of the optimum. Complementarily, [41] studies robustness of inexact VI using an invariant metric and proves convergence to a neighborhood of the optimum under sufficiently small disturbances; however, the result is still local, with bounds that depend on the initial value matrix. Finally, several model-free procedures guarantee only near-optimal convergence and may rely on partial model information. For example, [51] establishes near-optimal convergence for a specific LQR class (fast-sampling singular perturbed systems) using partial system knowledge, whereas [19] considers fully unknown systems but proves only convergence in probability and, as stated in Remark 5 in their article, requires the entire sequence of value iterates to be uniformly bounded, an assumption not guaranteed by the algorithm itself. The authors of [30, 29] study data-driven VI algorithms for systems subject to unmeasurable stochastic noise.
In this article, we study the convergence, robustness, and stability of VI for the stochastic LQR problem. First, we show that under noise-free conditions, VI exhibits global exponential convergence to the optimal solution from any positive semidefinite initial value function. Second, in the presence of noise in the VI procedure, we prove that VI converges to a small neighborhood of the optimal solution, provided that the noise is sufficiently small. Third, as a direct application of these results to adaptive dynamic programming (ADP) [31, 28], we develop a non-model-based robust least-squares value iteration (R-LSVI) algorithm for adaptive optimal controller design. We empirically demonstrate its convergence, stability, and adaptability and compare it with several established methods on a stochastic LQR problem – a stylized “data center cooling” scenario frequently used to benchmark reinforcement learning approaches [20, 58, 1, 18]. Finally, we also demonstrate our approach’s effectiveness on a dynamic portfolio allocation problem [21], a real-world problem from finance.
We emphasize that the convergence and robustness results presented in this article substantially relax the restrictive conditions on the initial value function and noise that are typically assumed in the existing literature [52, 43, 61, 41]. A central technique in our analysis is the adoption of input-to-state stability (ISS) from modern control theory [53, 54], specifically focusing on its recent variant, “small-disturbance ISS” [47, 15]. This framework facilitates a comprehensive analysis of VI’s convergence and robustness in the presence of disturbances and errors.
The remainder of this article is organized as follows. In Section 2, we present the necessary preliminaries, formulate the problem setup and summarize our main results. Section 3 introduces the R-LSVI algorithm, a fully data-driven method for constructing estimates of the value function directly from input-state data. In Section 4, we present a comprehensive convergence analysis of VI. We establish the global exponential stability of the exact VI method. This result is key in demonstrating that inexact VI achieves small-disturbance ISS. The convergence of the R-LSVI algorithm follows from it being a special case of inexact VI. In Section 5, we numerically demonstrate the convergence, stability, and adaptability of R-LSVI, comparing it to existing methods using a stochastic LQR problem inspired by a “data center cooling” scenario. In Section 6, we apply R-LSVI to a dynamic portfolio allocation problem, a practical application in finance. Section 7 concludes. We relegate some technical proofs to the appendix.
Notation: $\mathbb{R}$ (resp. $\mathbb{R}_{+}$) denotes the set of (nonnegative) real numbers; $\mathbb{Z}_{+}$ is the set of nonnegative integers; $\mathbb{S}^{n}$ is the set of real symmetric matrices of dimension $n$; $\mathbb{S}^{n}_{+}$ (resp. $\mathbb{S}^{n}_{++}$) is the set of real symmetric and positive semi-definite (definite) matrices of dimension $n$; $\otimes$ denotes the Kronecker product; $I_{n}$ denotes the identity matrix of dimension $n$ while $0_{m\times n}$ denotes the $m\times n$ zero matrix; $\|\cdot\|_{2}$ denotes the Euclidean norm for vectors and the spectral norm for matrices; for a function $\mathbb{F}\rightarrow\mathbb{R}^{n\times m}, u\|u\|_{\infty}$ denotes its $l^{\infty}$-norm when $\mathbb{F}=\mathbb{Z}_{+}$, and $L^{\infty}$-norm when $\mathbb{F}=\mathbb{R}$. $\mathbf{1}_{[0,T]}$ denotes an indicator function, that is $\mathbf{1}_{[0,T]}(t)=1$ if $t\in[0,T]$, and $\mathbf{1}_{[0,T]}(t)=0$ otherwise. For matrices $X\in\mathbb{R}^{m\times n},Y\in\mathbb{S}^{m}$, define $\mathrm{vec}(X)\triangleq\left[X_{1}^{\top},X_{2}^{\top},\ldots,X_{n}^{\top}\right]^{\top}\in\mathbb{R}^{nm}$, $\mathrm{svec}(Y)\triangleq\left[y_{11},\sqrt{2}y_{12},\ldots,\sqrt{2}y_{1m},y_{22},\sqrt{2}y_{23},\ldots,\sqrt{2}y_{m-1,m},y_{m,m}\right]^{\top}$ $\in\mathbb{R}^{\frac{1}{2}m(m+1)}$, where $X_{i}$ is the $i^{th}$ column of $X$. $\mathrm{vec}^{-1}(\cdot)$ and $\mathrm{svec}^{-1}(\cdot)$ are the operations such that $\mathrm{vec}^{-1}(\mathrm{vec}(X))=X$, and $\mathrm{svec}^{-1}(\mathrm{svec}(Y))=Y$. For $n\in\mathbb{Z}_{+}$, define $\tilde{n}\triangleq\frac{n(n+1)}{2}$. For a vector $v\in\mathbb{R}^{n}$, define $\tilde{v}\triangleq\mathrm{svec}(vv^{\top})\in\mathbb{R}^{\tilde{n}}$. For a vector $v\in\mathbb{R}^{n}$, define $\mathcal{E}(v)\triangleq\left[v_{1},\dots,v_{n-1}\right]^{\top}\in\mathbb{R}^{n-1}$ the truncation of the last element of $v$. For $Y\in\mathbb{S}^{n}$, $\lambda_{\mathrm{min}}\left(Y\right)$ ($\lambda_{\mathrm{max}}\left(Y\right)$) denotes its smallest (largest) algebraic eigenvalue. For $Y_{1},Y_{2}\in{\mathbb{S}^{n}}$, $Y_{1}\preceq Y_{2}$ means that $Y_{2}-Y_{1}\in\mathbb{S}^{n}_{+}$. For $S\in\mathbb{S}^{n}_{+}$, $S^{\frac{1}{2}}$ denotes the upper triangular Cholesky factor of $S$. For a matrix $\in\mathbb{R}^{n\times n}, ZZ^{\dagger}$ denotes its Moore-Penrose pseudoinverse. For $Z\in\mathbb{R}^{m\times n}$, define $\mathcal{B}_{r}(Z)\triangleq\left\{X\in\mathbb{R}^{m\times n},\lVert X-Z\rVert_{{2}}<r\right\}$ to be the open ball of radius $r$ centered at $Z$, and $\overline{\mathcal{B}}_{r}(Z)$ its closure, similarly $\mathcal{B}^{\infty}_{r}(Z)$ defines the ball for the infinity-norm.
2. Preliminaries and Problem Formulation
2.1. Exponential and Input-to-State Stability
In this article, we use exponential stability and input-to-state stability to analyze the convergence of VI. Below we recall the key definitions (see [54, 15] for details).
We consider the nonlinear system
$P_{i+1}=f(P_{i},\Delta_{i})$ (1)
where $P_{i}\in\mathbb{R}^{n}$ is the system state, $\Delta_{i}\in\mathbb{R}^{m}$ is the perturbation, $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is continuous, and $P^{*}$ is an equilibrium of $P_{i+1}=f\left(P_{i},0\right)$ when $\Delta_{i}=0$ for all $i\in\mathbb{Z}_{+}$.
Definition 1.
For the system (1) with $\Delta_{i}=0$ for all $i\in\mathbb{Z}_{+}$, an equilibrium $P^{*}$ is locally exponentially stable if there exists a $\delta>0$, such that for some $a>0$ and $0<b<1$,
$\lVert P_{i}-P^{*}\rVert_{2}\leq ab^{t}\lVert P_{0}-P^{*}\rVert_{2}$ (2)
for all $P_{0}\in\mathcal{B}_{\delta}\left(P^{*}\right)$. If $\delta=\infty$, then $P^{*}$ is a globally exponentially stable equilibrium.
Definition 2.
A function $\gamma:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ is said to be of class $\mathcal{K}$ if it is continuous, strictly increasing and vanishes at the origin. A function $\beta:\mathbb{R}_{+}\times\mathbb{Z}_{+}\rightarrow\mathbb{R}_{+}$ is said to be of class $\mathcal{KL}$ if $\beta(\cdot,i)$ is of class $\mathcal{K}$ for every fixed $i>0$, and for every fixed $r\geq 0$, $\beta(r,i)$ decreases to 0 as $i\rightarrow\infty$.
Definition 3.
The system (1) is small-disturbance input-to-state stable (ISS) if there exist some $d>0$, $\beta\in\mathcal{KL}$ and $\gamma\in\mathcal{K}$, such that for all $\Delta$ bounded by $d$ (i.e. $\left\|\Delta\right\|_{\infty}<d$) and all initial states $P_{0}\in\mathbb{R}^{n}$, $P_{i}$ satisfies
$\lVert P_{i}-P^{*}\rVert_{2}\leq\beta\left(\lVert P_{0}-P^{*}\rVert_{2},i\right)+\gamma\left(\lVert\Delta\rVert_{\infty}\right)\,.$ (3)
Exponential stability implies not only convergence of system (1), but also characterizes its rate of convergence. When the input signal is non-zero, the input-to-state stability describes how the solution of (1) is affected by the input signal. In particular, small-disturbance ISS ensures that the distance between the system state and the equilibrium remains uniformly bounded whenever the disturbance is sufficiently small, regardless of the initial state.
2.2. The Stochastic Linear Quadratic Regulator Problem in Discrete-Time
We consider a stochastic discrete-time linear system of the form
$x_{t+1}=Ax_{t}+Bu_{t}+C\epsilon_{t+1}$ (4)
where $x_{t}\in\mathbb{R}^{n}$ is the system state, $u_{t}\in\mathbb{R}^{m}$ is the control input, $\epsilon_{t+1}\in\mathbb{R}^{p}$ is an additive disturbance, $A\in\mathbb{R}^{n\times n},B\in\mathbb{R}^{n\times m}$ and $C\in\mathbb{R}^{n\times p}$. We assume that the initial state $x_{0}\in\mathbb{R}^{n}$ is given, $\mathbb{E}[\epsilon_{t}]=0_{p}$, $\mathbb{V}[\epsilon_{t}]=I_{p}$, and $\epsilon_{t},\epsilon_{t^{\prime}}$ are independent for all $t^{\prime}\neq t$. The stochastic LQR problem seeks a controller $u$ that minimizes the cost
$J\left(u\right)=\lim_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T-1}c\left(x_{t},u_{t}\right)\right]$ (5)
where $c(x_{t},u_{t})\triangleq x_{t}^{\top}Sx_{t}+u_{t}^{\top}Ru_{t}$ with $S\in\mathbb{R}^{n\times n}$ and $R\in\mathbb{R}^{m\times m}$.
Assumption 1.
$(A,B)$ is stabilizable, and $S\in\mathbb{S}^{n}_{++}$, $R\in\mathbb{S}^{m}_{++}$.
Under Assumption 1, the stochastic LQR problem admits a unique optimal controller $u^{*}_{t}=-K^{*}x_{t}$ with an optimal cost $J\left(u^{*}\right)=\mathrm{Tr}\left(C^{\top}P^{*}C\right)$, where $K^{*}\in\mathbb{R}^{m\times n}$ satisfies
$K^{*}=\left(R+B^{\top}P^{*}B\right)^{-1}B^{\top}P^{*}A$ (6)
and $P^{*}\in\mathbb{S}^{n}_{++}$ is the unique solution of the discrete algebraic Riccati equation (DARE)
$P$ $=A^{\top}PA-A^{\top}PB\left(R+B^{\top}PB\right)^{-1}B^{\top}PA+S\,.$ (7)
In addition, $A-BK^{*}$ is Schur stable, i.e. the spectral radius $\rho(A-BK^{*})<1$ (see, for example, [44, Section 2.4]). For any given $P\in\mathbb{S}_{+}^{n}$, the resulting control is
$K_{P}=\left(R+B^{\top}PB\right)^{-1}B^{\top}PA$ (8)
where the subscript in $K_{P}$ indicates its dependence on the value matrix $P$. For convenience, we write $K_{i}$ in place of $K_{P_{i}}$ throughout the article.
We introduce the Hamiltonian operator $\mathcal{Q}:\mathbb{R}^{n\times n}\to\mathbb{R}^{l\times l}$, with $l=n+m$,
$\mathcal{Q}\left(P\right)$ $\triangleq\left(\begin{array}[]{cc}A^{\top}PA+S&A^{\top}PB\\ B^{\top}PA&B^{\top}PB+R\end{array}\right)=\left(\begin{array}[]{cc}\left[\mathcal{Q}(P)\right]_{xx}&\left[\mathcal{Q}(P)\right]_{ux}^{\top}\\ \left[\mathcal{Q}(P)\right]_{ux}&\left[\mathcal{Q}(P)\right]_{uu}\\ \end{array}\right)$ (13)
which plays a fundamental role in the stochastic LQR problem. We define by $\mathcal{H}$ the $\cdot]_{uu}-$Schur complement operator, given by
$\mathcal{H}\left(Q\right)$ $\triangleq\left[{Q}\right]_{xx}-\left[{Q}\right]_{ux}^{\top}\left[{Q}\right]_{uu}^{-1}\left[{Q}\right]_{ux}\in\mathbb{R}^{n\times n}$ (14)
for any $Q\in\mathbb{R}^{l\times l}$ where $\left[{Q}\right]_{uu}$ is invertible. With this notation, we express the control gain and value matrix as
$K_{P}$ $=\left[\mathcal{Q}\left(P\right)\right]_{uu}^{-1}\left[\mathcal{Q}\left(P\right)\right]_{ux}$ (15)
$P$ $=\mathcal{H}\left(\mathcal{Q}(P)\right)\,.$ (16)
Since $P\in\mathbb{S}_{+}^{n}$ and $R\in\mathbb{S}_{++}^{m}$, it follows that $\left[\mathcal{Q}\left(P\right)\right]_{uu}\in\mathbb{S}_{++}^{m}$ and is therefore invertible. The formulas (15)–(16) capture the core principle behind the derivation of the Riccati equation. By eliminating the control variables via the Schur complement $\mathcal{H}$, we obtain an effective state cost that determines the optimal closed-loop behavior of the system.
To solve the stochastic LQR problem (4)–(5) we must solve the DARE (7) or equivalently (16). When the system dynamics are known, this can be accomplished using Bertsekas’s exact value iteration [7], as outlined in Procedure 1.
Procedure 1 (Exact Value Iteration).
  1. (1)
    Choose an initial value matrix $P_{0}\in\mathbb{S}^{n}_{+}$, and let $i=0$.
  2. (2)
    Update the value matrix via
    $P_{i+1}=\mathcal{H}\left(Q_{i}\right)$ (17)
    where $Q_{i}\triangleq\mathcal{Q}\left(P_{i}\right)$.
  3. (3)
    Set $i\leftarrow i+1$ and go back to Step (2).
We note that at any step of Procedure 1, the control gain $K_{i}$ associated with the value matrix $P_{i}$ can be explicitly computed as
$K_{i}=\left[\mathcal{Q}\left(P_{i}\right)\right]_{uu}^{-1}\left[\mathcal{Q}\left(P_{i}\right)\right]_{ux}.$ (18)
Unlike Hewer’s policy iteration [23], Procedure 1 converges even if $K_{i}$ is not stabilizing, as formally stated in the following theorem. This follows directly from [7, Proposition 4.4.1].
Theorem 1 (Convergence of Exact VI).
Under Assumption 1, for Procedure 1 it holds
  1. (i)
    $\lim_{i\rightarrow\infty}P_{i}=P^{*}$.
  2. (ii)
    $\lim_{i\rightarrow\infty}K_{i}=K^{*}$.
2.3. Problem Formulation
Theorem 1 states that exact VI (see Procedure 1) finds a sequence of approximations that are guaranteed to converge to the true solution $P^{*}$ and $K^{*}$. However, implementing exact VI requires evaluating $\mathcal{Q}$ defined in (13), which requires precise knowledge of $A$, $B$, $S$ and $R$.
In the setting of learning-based control, system dynamics are unknown and must be approximated using data-driven approaches. In some situations, the cost function may also be partially or completely unknown. Moreover, data employed in these approaches can be affected by disturbances such as system noise and numerical errors. Consequently, each iteration of the VI yields an approximate value matrix $\hat{P}_{i+1}$. To incorporate errors and noise, we formulate the following inexact version of the VI.
Procedure 2 (Inexact Value Iteration).
  1. (1)
    Choose an initial value matrix $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$, and let $i=0$.
  2. (2)
    Update the value matrix via
    $\hat{P}_{i+1}$ $=\mathcal{H}(\mathcal{Q}(\hat{P}_{i}))+\Delta_{i}\,.$
  3. (3)
    Set $i\leftarrow i+1$ and go back to Step (2).
In Procedure 2, $\Delta_{i}\in\mathbb{S}^{n}$ denotes external disturbances, while $\hat{P}_{i+1}$ is used to distinguish the sequence generated by the inexact value iteration from that of the exact value iteration. We always enforce the condition $\Delta_{i}\in\mathbb{S}^{n}$ by symmetrizing $\hat{P}_{i+1}=\frac{1}{2}(\hat{P}_{i+1}+\hat{P}_{i+1}^{\top})$.
Remark 1.
The inexact VI is general in the sense that any VI for stochastic LQR subject to errors or disturbances can be expressed as in Procedure 2, allowing us to analyze and establish conditions for its convergence, robustness and stability for a broad range of such perturbations. We provide two examples to illustrate this point.
[…]
$\Delta_{i}=\mathcal{H}(\hat{\mathcal{Q}}(\hat{P}_{i}))-\mathcal{H}({\mathcal{Q}}(\hat{P}_{i}))$ (19)
where
[…]
For the second, consider the R-LSVI algorithm (see Algorithm 1, below) where the source of disturbances arises from approximating the matrix $\mathcal{Q}(\hat{P}_{i})$. Then $\Delta_{i}$ takes the form
$\Delta_{i}=\mathcal{H}(\hat{\mathcal{Q}}(\hat{P}_{i}))-\mathcal{H}(\mathcal{Q}(\hat{P}_{i}))$ (23)
[…]
$\hat{\mathcal{Q}}\left(P\right)$ $\triangleq\mathrm{svec}^{-1}(\mathcal{E}(\Theta_{T}^{\dagger}\Psi_{T}\mathrm{svec}(P)+\Theta_{T}^{\dagger}\Xi_{T}))\,.$ (24)
In both examples, the error arises from approximating the Hamiltonian operator $\mathcal{Q}$, but the nature of the approximation differs significantly between the two. In the first example, known as the nominal control approach, the system and cost matrices $(A,B,S,R)$ are approximated first, and then the Hamiltonian is constructed in a model-based manner. In contrast, in the second example, based on adaptive dynamic programming (ADP), we directly approximate the Hamiltonian, resulting in a non-model-based approach. This distinction affects the structure of the perturbations and the resulting error propagation in the inexact value iteration process. In this article, we establish sufficient conditions on $\{\Delta_{i}\}_{i=0}^{\infty}$ to guarantee the convergence of sequences generated by inexact VI to a neighborhood of the true solution $P^{*}$.
2.4. Main Contributions
Our contributions are threefold:
Result 1: We show that the exact VI (see Procedure 1) is globally exponentially stable for any $P_{0}\in\mathbb{S}^{n}_{+}$.
Result 2: We prove that the inexact VI (see Procedure 2) is small-disturbance ISS.
Result 3: We propose a new non-model-based adaptive dynamic programming algorithm based on VI to determine the optimal controller in (6) using only input-state data, with convergence guarantees.
In addition, we evaluate empirically the convergence, stability, and adaptability of our proposed algorithm using the “data center cooling” stochastic LQR problem, a benchmark frequently employed to assess various RL approaches (e.g. [20, 58, 1, 18]). Our simulations compare the proposed algorithm with several alternatives, including both model-based PI and VI approaches [50, 18], a model-free policy gradient method [56], and two ADP policy iteration procedures [40, 47]. The results highlight the superior adaptability and stability of our approach, even when costs are non-quadratic.
3. Robust Least-Squares VI
We introduce the robust least-squares VI (R-LSVI) algorithm in this section, a fully data-driven approach to estimate $\hat{P}_{i}$ in Procedure 2 from input-state data. Section 4.3 provides a rigorous convergence analysis.
We define the filtration $\{\mathcal{F}_{t}\}_{t\geq 0}$ by $\mathcal{F}_{t}=\sigma(x_{0},\ldots,\allowbreak x_{t})$, where $\sigma(\cdot)$ denotes the $\sigma$-algebra generated by the state sequence of up to time $t$.
Because $(A,B,S,R)$ are unknown, the behavior policy $u_{t}$ we choose for data collection may produce an unstable closed loop and drive the state to infinity. To prevent divergence and numerical ill-conditioning, we reset the state to $0_{n}$ whenever $\lVert x_{t}\rVert_{\infty}$ exceeds $d$. We formalize this mechanism in the following proposition.
Proposition 1.
Define the reset dynamics with initial condition $x_{0}=0_{n}$
$X_{t+1}$ $=Ax_{t}+Bu_{t}+C\epsilon_{t+1},\quad\epsilon_{t}\underset{iid}{\sim}\mathcal{N}(0_{p},I_{p})$ (25)
$x_{t+1}$ $=\mathbbm{1}_{{\lVert X_{t}\rVert_{\infty}\leq d}}X_{t+1},\quad d>0$ (26)
and consider a behavior policy of the form
$u_{t}=-K_{c}x_{t}+\eta_{t+1},\quad\eta_{t}\underset{iid}{\sim}\mathcal{N}(0_{m},\Sigma_{\eta}),\quad\Sigma_{\eta}\succ 0$ (27)
where $K_{c}$ is not necessarily stabilizing. Then the Markov chain $(x_{t})_{t\geq 0}$ admits a unique invariant probability measure $\pi$.
The proof relies on the irreducibility and positive recurrence of $(x_{t})_{t\geq 0}$ (via regeneration at $0_{n}$), which ensure the existence and uniqueness of the invariant measure $\pi$; see Appendix B for details.
We now derive the least squares formulation used in the construction of $\hat{\mathcal{Q}}(P)$. For any P $\in\mathbb{S}^{n}_{+}$, recall $X_{t+1}=Ax_{t}+Bu_{t}+C\epsilon_{t+1}$, and hence
$\mathbb{E}\left[X_{t+1}^{\top}PX_{t+1}+c_{t}|\mathcal{F}_{t}\right]$ (28)
$\quad=(Ax_{t}+Bu_{t})^{\top}P(Ax_{t}+Bu_{t})+x_{t}^{\top}Sx_{t}+u_{t}^{\top}Ru_{t}$ (29)
$\quad\qquad+\mathbb{E}\left[(C\epsilon_{t+1})^{\top}P(C\epsilon_{t+1})|\mathcal{F}_{t}\right]$ (30)
$\quad=\left(\begin{array}[]{c}x_{t}\\ u_{t}\end{array}\right)^{\top}\mathcal{Q}(P)\left(\begin{array}[]{c}x_{t}\\ u_{t}\end{array}\right)+\mu(P)$ (35)
where $\mu(P)=\mathrm{Tr}\left(C^{\top}PC\right)$. Setting $y_{t}=[x_{t}^{\top},u_{t}^{\top}]^{\top}$, we obtain
$\mathbb{E}\left[\tilde{X}_{t+1}^{\top}\mathrm{svec}(P)|\mathcal{F}_{t}\right]+c_{t}=\tilde{y}_{t}^{\top}\mathrm{svec}(\mathcal{Q}\left(P\right))+\mu\left(P\right)\,.$ (36)
Defining $z_{t}\triangleq\left[\tilde{y}_{t}^{\top},1\right]^{\top}$, multiplying both sides of (36) by $\frac{z_{t}}{\alpha_{t}^{2}}$ with $\alpha_{t}=\lVert z_{t}\rVert_{\infty}$, and taking the expectation with respect to the invariant distribution $\pi$ yields
$\Theta\left(\begin{array}[]{l}\mathrm{svec}(\mathcal{Q}\left(P\right))\\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mu\left(P\right)\end{array}\right)=\Psi\mathrm{svec}(P)+\Xi$ (37)
where
$\Theta\triangleq\mathbb{E}_{\pi}\left[\frac{z_{t}z_{t}^{\top}}{\alpha_{t}^{2}}\right],\quad\Psi\triangleq\mathbb{E}_{\pi}\left[\frac{z_{t}\tilde{X}_{t+1}^{\top}}{\alpha_{t}^{2}}\right],\quad\Xi\triangleq\mathbb{E}_{\pi}\left[\frac{z_{t}c_{t}}{\alpha_{t}^{2}}\right].$ (38)
Assumption 2.
There exists $c>0$, such that for $\phi_{t}\triangleq\frac{z_{t}}{\alpha_{t}}$ we have
$\Theta=\mathbb{E}_{\pi}[\phi_{t}\phi_{t}^{\top}]\succeq cI_{\tilde{l}+1}\,.$ (39)
Remark 2.
Assumption 2 is consistent with the persistent excitation condition widely used in adaptive control [4, 60] and is similarly prevalent in various RL methods (see, for example, [28], [34], [36]).
For a given $P$ on the right hand side of (37), we can estimate $\mathcal{Q}\left(P\right)$ on the left hand side from collected data via least-squares (LS), leading to
$\left(\begin{array}[]{l}\mathrm{svec}(\hat{\mathcal{Q}}\left(P\right))\\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \hat{\mu}\left(P\right)\end{array}\right)=\Theta_{T}^{\dagger}\Psi_{T}\mathrm{svec}(P)+\Theta_{T}^{\dagger}\Xi_{T}$ (40)
where the data matrices
$\Theta_{T}\triangleq\frac{1}{T}\sum_{t=1}^{T}\frac{z_{t}z_{t}^{\top}}{\alpha_{t}^{2}},\penalty 10000\ \Psi_{T}\triangleq\frac{1}{T}\sum_{t=1}^{T}\frac{z_{t}\tilde{X}_{t+1}^{\top}}{\alpha_{t}^{2}},\penalty 10000\ \Xi_{T}\triangleq\frac{1}{T}\sum_{t=0}^{T-1}\frac{z_{t}c_{t}}{\alpha_{t}^{2}}$ (41)
are sample estimators of (38). As the chain $(x_{t})_{t>0}$ is irreducible and positive recurrent, by the ergodic theorem, the data matrices in (41) converge with increasing data size, i.e.
$\lim_{T\rightarrow\infty}\Theta_{T}=\Theta,\quad\lim_{T\rightarrow\infty}\Psi_{T}=\Psi,\quad\lim_{T\rightarrow\infty}\Xi_{T}=\Xi\,.$ (42)
Remark 3.
The rescaling by $\alpha_{t}=\lVert z_{t}\rVert_{\infty}$ in (38) and (41) acts as a numerical preconditioning step. Because $\lVert x_{t}\rVert_{\infty}\in[0,d]$ (with $d$ potentially large), the dynamic range of the data can be wide. Without normalization, large-magnitude samples dominate the averages and yield ill-conditioned estimates of (41).
We now present the R-LSVI algorithm in Algorithm 1. The method operates entirely without access to system matrices $(A,B,S,R)$. We sample data by applying the behavior policy $u_{t}$ defined in (27) and observing $(X_{t},x_{t},c_{t})$ along trajectories of (25),(26) and (5), from which we construct the data matrices (41). The subsequent steps follow Procedure 1, with $\hat{Q}_{i}$, computed via (40), replacing $Q_{i}$. Finally, we obtain the control gain $\hat{K}$ using formula (15), as it is the primary quantity of practical interest. Note that contrary to the PI Procedure, we do not need the knowledge of a stabilizing initial controller $K_{0}$.
Algorithm 1 Robust Least-Squares Value Iteration (R-LSVI)
input: Choose $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$, exploration-noise covariance $\Sigma_{\eta}$, number of iterations $I_{\text{max}}$, time horizon $T$, the reset bound $d>0$, and let $i=0$.
 Collect data $(X_{t},x_{t},u_{t})$ from (25),(26) and (27).
 Construct $\Theta_{T}$, $\Psi_{T}$ and $\Xi_{T}$.
while $i<I_{\text{max}}$ do
  $\hat{Q}_{i}=\mathrm{svec}^{-1}(\mathcal{E}(\Theta_{T}^{\dagger}\Psi_{T}\mathrm{svec}(\hat{P}_{i})+\Theta_{T}^{\dagger}\Xi_{T}))$
  $\hat{P}_{i+1}=\mathcal{H}(\hat{Q}_{i})$
  Update $i\leftarrow i+1$.
end while
output: $\hat{K}=[\hat{Q}_{I_{\text{max}}}]_{uu}^{\dagger}[\hat{Q}_{I_{\text{max}}}]_{ux}$
Remark 4.
This procedure significantly sharpens previous results. Specifically, the law of $(x_{t})$ converges to the invariant distribution $\pi$ leading to almost sure convergence of the data matrices, whereas prior work established only convergence in probability (e.g. [19]).
Remark 5.
We make the following remarks regarding the data matrices (41) in the R-LSVI algorithm:
  1. (i)
    The algorithm computes $\Theta_{T}^{\dagger}\Psi_{T}$ and $\Theta_{T}^{\dagger}\Xi_{T}$ once and reuses them in every iteration, which improves computational efficiency. The overall complexity is dominated by computing the pseudo inverse of $\Theta_{T}$ and the matrix products performed at each step. The overall complexity is $O(T\tilde{l}^{2}+I_{\max}\tilde{l}^{2})$.
  2. (ii)
    R-LSVI constructs the data matrices entirely from input-state data and, unlike the approaches in [13, 27, 31, 9], does not rely on observing system disturbances $\{\epsilon_{t}\}_{t\geq 0}$.
  3. (iii)
    The data matrices are smaller in size than the ones typically found in the ADP literature (see, for example, [26], [48]). For instance, $\Theta_{T}$ is of size $(\tilde{l}+1)^{2}$ rather than $(\tilde{l}+l+1)^{2}$ . This dimension reduction arises from considering $\left(\begin{array}[]{l}\mathrm{svec}(\mathcal{Q}\left(P\right))\\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \mu\left(P\right)\end{array}\right)$ rather than $\left(\begin{array}[]{ll}\mathrm{svec}(\mathcal{Q}\left(P\right))&0_{(n+m)}\\ 0_{(n+m)}^{\top}&\penalty 10000\ \penalty 10000\ \penalty 10000\ \mu\left(P\right)\end{array}\right)$ in the LS problem (37), leading to a mathematically equivalent but computationally more efficient formulation.
4. Convergence Analysis
In the inexact VI defined in Procedure 2, the updates take the form
$\hat{P}_{i+1}$ $=\mathcal{H}(\mathcal{Q}(\hat{P}_{i}))+\Delta_{i}$ (43)
$=A^{\top}\hat{P}_{i}A-A^{\top}\hat{P}_{i}B\left(R+B^{\top}\hat{P}_{i}B\right)^{-1}B^{\top}\hat{P}_{i}A+S+\Delta_{i}\,.$ (44)
We observe that (44) can be written as
$\hat{P}_{i+1}$ $=\mathcal{R}(\hat{P}_{i},S)+\Delta_{i}=\mathcal{R}(\hat{P}_{i},S+\Delta_{i})$ (45)
where $\mathcal{R}$ is the Riccati operator defined by
$\mathcal{R}\left(P,S\right)\triangleq A^{\top}PA-A^{\top}PB(R+B^{\top}PB)^{-1}B^{\top}PA+S$ (46)
for $P\in\mathbb{S}^{n}_{+}$. The operator $\mathcal{R}$ plays a central role in our analysis. Its structure allows us to reduce the study of $\hat{P}_{i+1}$’s dependence on $\Delta_{i}$ to understanding how the Riccati solution depends on the cost matrix $S$.
4.1. Convergence of the Exact VI
We begin by establishing the convergence of exact VI in this section, laying the groundwork for proving the convergence and robustness of inexact VI. Detailed proofs can be found in Appendix C.
The following theorem shows that exact VI is globally exponentially stable, thereby extending earlier results, which either establish local exponential stability [52] or guarantee exponential stability only when the initial value function lies within the semidefinite cone of the optimal solution [43, 61].
Theorem 2 (Global Exponential Convergence of Exact VI).
The exact value iteration in Procedure 1 is globally exponentially stable. That is, there exists a $\mathcal{KL}$-function $\beta$, such that $\forall i\in\mathbb{Z}_{+}\text{ and }\forall P_{0}\in\mathbb{S}_{+}^{n}$
$\lVert P_{i}-P^{*}\rVert_{2}\leq\beta(\lVert P_{0}-P^{*}\rVert_{2},i).$ (47)
In addition, $\beta(r,i)={\alpha}\lVert A-BK^{*}\rVert_{2}^{i}r$ for some constant ${\alpha}\geq 1$.
The next proposition asserts that a sequence of value matrices generated by iterating on the DARE will dominate another such sequence (in the $\succeq$ sense), if the cost matrices associated with the first sequence dominate those of the second. This result is the discrete analogue of applying Grönwall’s inequality [57] to the matrix differential equation $\dot{P}(t)=\mathcal{R}(P(t),S(t))$.
Proposition 2.
Let $\{P_{i}^{j}\}_{i\in\mathbb{Z}_{+}}$ with $j\in\{1,2\}$ be the following sequences
$P_{i+1}^{j}$ $=\mathcal{R}(P_{i}^{j},S_{i}^{j}),\quad P_{0}^{j}\in\mathbb{S}^{n}_{+}.$ (48)
Suppose that $S^{1}_{i}\succeq S^{2}_{i}$ for all $i\in\mathbb{Z}_{+}$ and $P_{0}^{1}\succeq P_{0}^{2}$. Then, $P_{i}^{1}\succeq P_{i}^{2}$ for all $i\in\mathbb{Z}_{+}$.
Proposition 3 states that the unique solution to the Riccati equation with perturbed cost matrix $S+\delta I_{n}$ depends smoothly on the perturbation parameter $\delta$.
Proposition 3.
Let $\delta>-\lambda_{\mathrm{min}}\left(S\right)$ and $P^{*}(\delta)\in\mathbb{S}^{n}_{++}$ be the unique solution of the following DARE:
$P$ $=\mathcal{R}\left(P,S+\delta I_{n}\right).$ (49)
Then $P^{*}(\delta)$ is continuously differentiable for all $\delta\in(-\lambda_{\mathrm{min}}\left(S\right),\infty)$.
4.2. Convergence of the Inexact VI
This section focuses on the analysis of inexact VI in Procedure 2, which serves as a general framework for examining the convergence properties of VI for a broad range of disturbances. Using Proposition 2, we first show that inexact VI can be bounded between two instances of exact VI. Then, leveraging Proposition 3, we demonstrate that these bounds can be made arbitrarily tight in a continuous manner, concluding that inexact VI is small-disturbance ISS.
Theorem 3 (Small-Disturbance ISS of Inexact VI).
Inexact value iteration is small-disturbance ISS. Specifically, if $\lVert\Delta\rVert_{\infty}<\lambda_{\mathrm{min}}\left(S\right)$, there exist a $\mathcal{KL}$-function $\beta$, and a $\mathcal{K}$-function $\gamma$, such that
$\lVert\hat{P}_{i}-P^{*}\rVert_{2}\leq\beta(\lVert\hat{P}_{0}-P^{*}\rVert_{2},i)+\gamma(\lVert\Delta\rVert_{\infty})$ (50)
for all $i\in\mathbb{Z}_{+}$ and all $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$.
Proof.
Define $\delta^{+}\triangleq\lVert\Delta\rVert_{\infty}$, $\delta^{-}\triangleq-\lVert\Delta\rVert_{\infty}$, and consider the sequences $\{P_{i}^{+}\}_{i\in\mathbb{Z}+}$ and $\{P_{i}^{-}\}_{i\in\mathbb{Z}+}$ generated by
${P}_{i+1}^{+}$ $=\mathcal{R}\left({P}_{i}^{+},S+\delta^{+}I_{n}\right),\quad{P}^{+}_{0}=P_{0}$ (51)
${P}_{i+1}^{-}$ $=\mathcal{R}\left({P}_{i}^{-},S+\delta^{-}I_{n}\right),\quad{P}^{-}_{0}=P_{0}.$ (52)
Since the pair $(A,B)$ is stabilizable, $S+\delta^{+}I_{n}\succ 0$ and $S+\delta^{-}I_{n}\succ 0$, then both ${P}^{+}_{i}$ and ${P}^{-}_{i}$ converge to ${P}_{+}^{*}$ and ${P}_{-}^{*}$, which are the unique positive-definite solutions of
${P}_{+}^{*}$ $=\mathcal{R}\left({P}_{+}^{*},S+\delta^{+}I_{n}\right)$ (53)
${P}_{-}^{*}$ $=\mathcal{R}\left({P}_{-}^{*},S+\delta^{-}I_{n}\right).$ (54)
By Proposition 3, there exists an $L>0$ such that $\lVert P_{+}^{*}-P^{*}\rVert_{2}\leq L\lVert\Delta\rVert_{\infty}$ and $\lVert P_{-}^{*}-P^{*}\rVert_{2}\leq L\lVert\Delta\rVert_{\infty}$. By Theorem 2 and the weak triangle inequality [32] we obtain
$\lVert P_{i}^{+}-P_{+}^{*}\rVert_{2}$ $\leq\beta(\lVert P_{0}-P_{+}^{*}\rVert_{2},i)\leq\beta(\lVert P_{0}-P^{*}\rVert_{2},i)+\gamma(\lVert\Delta\rVert_{\infty})$ (55)
$\lVert P_{i}^{-}-P_{-}^{*}\rVert_{2}$ $\leq\beta(\lVert P_{0}-P_{-}^{*}\rVert_{2},i)\leq\beta(\lVert P_{0}-P^{*}\rVert_{2},i)+\gamma(\lVert\Delta\rVert_{\infty}).$ (56)
By Proposition 2 we have
$P_{i}^{-}-P^{*}\preceq\hat{P}_{i}-P^{*}\preceq P_{i}^{+}-P^{*}.$ (57)
Hence, we have
$\lVert\hat{P}_{i}-P^{*}\rVert_{2}$ $\leq\max\{\lVert P_{i}^{+}-P^{*}\rVert_{2},\lVert P_{i}^{-}-P^{*}\rVert_{2}\}$
$\leq\beta(\lVert P_{0}-P^{*}\rVert_{2},i)+\gamma(\lVert\Delta\rVert_{\infty})$ (58)
and the proof is complete. ∎
Corollary 1.
For a given $\hat{P}_{0}\in\mathbb{S}_{+}^{n}$, let $\mathcal{C}(\hat{P}_{0})$ denote the set of value matrices generated by the inexact value iteration in Procedure 2, starting from $\hat{P}_{0}$, with all possible $\left\{\Delta_{i}\right\}_{i=0}^{\infty}$ satisfying $\|\Delta\|_{\infty}<\lambda_{\mathrm{min}}\left(S\right)$. Then the closure $\bar{\mathcal{C}}(\hat{P}_{0})$ is compact.
Remark 6.
Theorem 3 provides a global result that extends the local exponential asymptotic stability results of continuous-time systems shown in [11] and the exponential asymptotic stability for any $\hat{P}_{0}\succeq P^{*}$ established in [43, 61].
4.3. Convergence of R-LSVI
In this section, we establish the convergence of the R-LSVI algorithm by demonstrating that it is a special case of inexact VI. Disturbances in the R-LSVI algorithm (Algorithm 1) result from approximation errors in $Q_{i}=\mathcal{Q}(\hat{P}_{i})$ of the form $\hat{Q}_{i}=Q_{i}+\Delta Q_{i}$, which are due to insufficient data, system noise, and numerical inaccuracies. Theorem 4 establishes the convergence of R-LSVI.
Proposition 4.
For any $\hat{P}_{0}\in\mathbb{S}_{+}^{n}$, there exists $d^{*}>0$, a $\mathcal{KL}$-function $\beta$ and a $\mathcal{K}$-function $\tilde{\gamma}$, such that if $\lVert\Delta Q\rVert_{\infty}<d^{*}$, then
$\lVert\hat{P}_{i}-P^{*}\rVert_{2}\leq\beta(\lVert\hat{P}_{0}-P^{*}\rVert_{2},i)+\tilde{\gamma}(\lVert\Delta Q\rVert_{\infty})$ (59)
for all $i\in\mathbb{Z}_{+}$.
Proof.
Consider the set $\mathcal{C}(\hat{P}_{0})$ of $\mathbb{S}^{n}_{+}$ defined in Corollary 1. By Lemma 3, we know that for $\lVert\Delta Q_{i}\rVert_{2}<d=\min(\lambda_{\mathrm{min}}\left(S\right),\lambda_{\mathrm{min}}\left(R\right))$, $Q_{i}]_{uu}$ and $\hat{Q}_{i}]_{uu}$ are invertible. Therefore, we relate $\Delta_{i}$ and $\Delta Q_{i}$ via
$\Delta_{i}$ $=\mathcal{H}(\hat{Q}_{i})-\mathcal{H}\left(Q_{i}\right)$ (60)
where $\mathcal{H}$ is an analytic function. The set $\{Q_{i}=\mathcal{Q}(\hat{P}_{i})|\hat{P}_{i}\in\bar{\mathcal{C}}(\hat{P}_{0})\}$ is compact as it is the image of a compact set under the continuous function $\mathcal{Q}$. By the extreme value theorem there exists a constant $c(\mathcal{C}(\hat{P}_{0}))>0$ such that
$\lVert\Delta_{i}\rVert_{\infty}\leq c(\mathcal{C}(\hat{P}_{0}))\lVert\Delta Q_{i}\rVert_{\infty}\,.$ (61)
Therefore, with $d^{*}\triangleq\min\left(d,\frac{\lambda_{\mathrm{min}}\left(S\right)}{2c(\mathcal{C}(\hat{P}_{0}))}\right)$, the sequence $\{\Delta_{i}\}_{i\geq 0}$ generated by Algorithm 1 satisfies
$\left\|\Delta_{i}\right\|_{\infty}<\lambda_{\mathrm{min}}\left(S\right)$ (62)
for all $i>0$. By Theorem 3, we obtain
$\lVert\hat{P}_{i}-P^{*}\rVert_{2}$ $\leq\beta(\lVert\hat{P}_{0}-P^{*}\rVert_{{2}},i)+\gamma(\lVert\Delta\rVert_{\infty})$
$\leq\beta(\lVert\hat{P}_{0}-P^{*}\rVert_{2},i)+\gamma(c(\mathcal{C}(\hat{P}_{0}))\lVert\Delta Q_{i}\rVert_{\infty})\,.$ (63)
To conclude the proof, set the $\mathcal{K}$-function to $\tilde{\gamma}(x)\triangleq\gamma(c(\mathcal{C}(\hat{P}_{0}))x)$. ∎
The following proposition shows that there exists $T_{0}>0$ independent of the iteration step $i$, such that for a sample size $T>T_{0}$, Algorithm 1 yields a $\Delta\lVert Q_{i}\rVert_{\infty}<d^{*}$.
Proposition 5.
For a given $\hat{P}_{0}$, let $\mathcal{P}(\hat{P}_{0})$ denote the set of value matrices generated by Algorithm 1, starting from $\hat{P}_{0}$, with all possible $\left\{\Delta Q_{i}\right\}_{i=0}^{\infty}$ satisfying $\|\Delta Q\|_{\infty}<d^{*}$ with $d^{*}$ defined in Theorem 4. Then, under Assumption 2 there exist $T_{0}\in\mathbb{Z}^{+}$, such that for any $T>T_{0}$
$\hat{P}_{i}\in\mathcal{P}(\hat{P}_{0})\implies\|\Delta Q_{i}\|_{2}<d^{*},\text{ almost surely}.$ (64)
Proof.
For all $\hat{P}_{i}\in\mathcal{P}(\hat{P}_{0})$, the sequence is bounded. Indeed by construction, the sequence $\{\Delta Q_{i}\}_{i\geq 0}$ induce $\{\Delta_{i}\}_{i\geq 0}$ such that $\lVert\Delta\rVert_{\infty}<d^{*}\leq\lambda_{\mathrm{min}}\left(S\right)$. Hence $\mathcal{P}(\hat{P}_{0})\subset\mathcal{C}(\hat{P}_{0})$, and by Corollary 1, $\bar{\mathcal{C}}(\hat{P}_{0})$ is a compact set. Therefore, $\bar{\mathcal{P}}(\hat{P}_{0})$ is also a compact set. From Algorithm 1 we have that
$\hat{Q}_{i}=\mathrm{svec}^{-1}(\mathcal{E}(\Theta^{\dagger}_{T}\Psi_{T}\mathrm{svec}(\hat{P}_{i})+\Theta^{{\dagger}}_{T}\Xi_{T}))\,.$ (65)
Let $\mathcal{T}$ be the operator defined by
$\hat{Q}_{i}=\mathcal{T}(\Theta_{T},\Psi_{T},\Xi_{T},\hat{P}_{i})\,.$ (66)
Recall that ${Q}_{i}=\mathcal{Q}(\hat{P}_{i})$ and $\Delta Q_{i}=\hat{Q}_{i}-{Q}_{i}$, and hence
$\Delta Q_{i}=\mathcal{T}(\Theta_{T},\Psi_{T},\Xi_{T},\hat{P}_{i})-\mathcal{T}(\Theta,\Psi,\Xi,\hat{P}_{i})$ (67)
where
$\mathcal{T}\left(\Theta,\Psi,\Xi,\hat{P}_{i}\right)=\left(\begin{array}[]{cc}A^{\top}\hat{P}_{i}A+S&A^{\top}\hat{P}_{i}B\\ B^{\top}\hat{P}_{i}A&B^{\top}\hat{P}_{i}B+R\end{array}\right)\,.$ (70)
By the ergodic theorem, we have
$\lim_{N\rightarrow\infty}\Theta_{T}=\Theta,\quad\lim_{N\rightarrow\infty}\Psi_{T}=\Psi,\quad\lim_{N\rightarrow\infty}\Xi_{T}=\Xi$ (71)
with $\Theta_{T}$ invertible by Assumption 2. By the continuity of matrix inverse theorem, there exists $\epsilon>0$ such that for all $\Theta_{T}\in\mathcal{B}_{\epsilon}\left(\Theta\right)$, $\Theta_{T}$ is invertible. Hence there exists $T_{1}$ such that for all $T>T_{1}$, $\Theta_{T}\in\mathcal{B}_{\epsilon}\left(\Theta\right)$. Furthermore, the function $\mathcal{T}(x,y,z,p)$ is continuous in $(x^{-1},y,z)$ and analytic in $p$, therefore, for any $\hat{P}_{i}\in\bar{\mathcal{P}}(\hat{P}_{0})$ and any $d^{*}>0$, there exists $T_{\hat{P}_{i}}>T_{1}$ such that for all $T>T_{\hat{P}_{i}}$
$\left\|\Delta Q_{i}\right\|_{2}$ $=\lVert\mathcal{T}(\Theta_{T},\Psi_{T},\Xi_{T},\hat{P}_{i})-\mathcal{T}(\Theta,\Psi,\Xi,\hat{P}_{i})\rVert_{2}<d^{*}\,.$ (72)
Finally, because $\bar{\mathcal{P}}(\hat{P}_{0})$ is compact, there exists $T_{0}\in\mathbb{Z}^{+}$ such that for all $T>T_{0}>T_{1}$, equation (72) is verified. ∎
We proceed to prove the convergence of R-LSVI.
Theorem 4 (Convergence of R-LSVI).
Under Assumption 2, for Algorithm 1, given any initial value matrix $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$ and any $\epsilon>0$, there exists $T_{0}\in\mathbb{Z}_{+}$ such that for all $T\geq T_{0}$
$\limsup_{i\rightarrow\infty}\lVert\hat{P}_{i}-P^{*}\rVert_{2}<\epsilon\text{ almost surely}\,.$
Proof.
Since $\hat{P}_{0}\in\mathcal{P}(\hat{P}_{0})$, by Proposition 5, $\left\|\Delta Q_{1}\right\|_{2}<d^{*}$ almost surely. By definition, $\hat{P}_{1}\in\mathcal{P}(\hat{P}_{0})$. Thus, by induction, we have $\left\|\Delta Q\right\|_{\infty}<d^{*}$ almost surely. By Theorem 4, the result follows. ∎
5. Empirical Results
We examine the convergence, stability, and robustness of the R-LSVI algorithm using a standard RL benchmark problem with the unstable Laplacian dynamics, given by (4)–(5) with
$A=\left[\begin{array}[]{lll}1.01&0.01&0\\ 0.01&1.01&0.01\\ 0&0.01&1.01\\ \end{array}\right],\quad B=\left[\begin{array}[]{lll}1&0&0\\ 0&1&0\\ 0&0&1\\ \end{array}\right]$ (73)
$C=S=I_{3},\quad R=1000I_{3}\,.$ (74)
Several authors have adopted this problem, which represents a stylized “data center cooling” scenario [20], to evaluate different RL approaches for stochastic LQR [58, 1, 18].
Remark 7.
The open-loop system given by the matrices in (73) and (74) is unstable. Specifically, for any nonzero initial condition, the open-loop state sequence diverges. Should an RL algorithm (explicitly or implicitly) underestimate a diagonal entry of $A$ as less than one, it will erroneously treat the corresponding mode as stable and allocate insufficient control effort, especially given the high control cost represented by $R=1000I_{3}$. Consequently, obtaining a high-quality estimate of the system’s true behavior is imperative for achieving near-optimal control.
For comparison, we include the following well-established methods:
  • Nominal Control [50, 18]: a plugin PI (VI) model-based procedure that first estimates the system and cost matrices $(A,B,Q,R)$ and then solves the stochastic LQR problem using Hewer’s PI [23] (Bertsekas’s VI see [7]).
  • Policy Gradient [56]: a model-free approach that directly learns a policy function from episodic experiences. In our implementation, we use the Adam optimizer [35] for policy gradient updates and estimate the cost with the running cost over ten consecutive points. We supply the algorithm with the exact gradient.
  • LSPI [40]: a classical PI-based ADP method. For a fair comparison with R-LSVI, we adapt LSPI to incorporate the additive noise term $\mu(P)$ and to include the cost matrices $(S,R)$ explicitly in the Hamiltonian $\mathcal{Q}(P)$.
  • O-LSPI [47]: an ADP procedure consisting of two loops, an outer PI loop and an inner VI loop. We set $I_{\text{max}}^{\text{inner}}=20$ and $I_{\text{max}}^{\text{outer}}=5$ to enhance the overall accuracy of the policy updates.
[…]
Remark 8.
Each methods yields a control gain $\hat{K}$. If $\hat{K}$ is stabilizing, the linear state feedback control policy $u_{t}=-\hat{K}x_{t}$ is applied to the system, resulting the cost $J\left(u\right)=\mathrm{Tr}\left(C^{\top}\hat{P}C\right)$, where the value matrix $\hat{P}\in\mathbb{S}^{n}_{++}$ is the unique positive definite solution of the Lyapunov equation
$\hat{P}=(A-B\hat{K})^{\top}\hat{P}(A-B\hat{K})+S+\hat{K}^{\top}R\hat{K}\,$. (75)
This distinction matters for value iteration because the controller is derived from $\hat{P}$ (computed via the Lyapunov equation (75)) rather than from $\hat{P}_{{I_{\text{max}}}}$.
For each problem instance and method, we run the algorithms $100$ times and calculate the fraction of trials that produce a stabilizing control gain $\hat{K}$. To ensure comparability, all algorithms are executed for a total of $100$ iterations $\left(I_{\text{max}}=100\right)$ and use the same number of data points.
We initialize each procedure with $x_{\text{init}}=0_{3}$ and $\hat{P}_{0}=\beta I_{3}$, with $\beta\sim\mathcal{U}\left(0,1\right)$ to ensure that $\hat{P}_{0}\in\mathbb{S}^{n}_{+}$. Since the nominal PI, LSPI and O-LSPI are PI-based algorithms, it is necessary to select an initial control gain, $\hat{K}_{0}$. For consistency and fairness across comparisons, we choose $\hat{K}_{0}$ as the corresponding induced gains $\hat{K}_{0}=(R+B^{\top}\hat{P}_{0}B)^{-1}B^{\top}\hat{P}_{0}A$, as in [52]. For the policy gradient specifically, we take $\hat{K}_{0}$ stabilizing, $\hat{K}_{0}=\alpha_{0}I_{3}$, with $\alpha_{0}\sim\mathcal{U}\left(0.1,1\right)$ stabilizing to enhance its convergence.
5.1. Convergence Analysis
We evaluate algorithmic convergence on the benchmark problem as the sample size increases In order to address the non-stabilizing collect policy. In our experiments, we assess convergence using the cost relative error, defined as
$\frac{\mathrm{Tr}\left(C^{\top}(\hat{P}-P^{*})C\right)}{\mathrm{Tr}\left(C^{\top}P^{*}C\right)}$ (76)
where the optimal $P^{*}$ is given by formula (7). In addition, we compute the fraction of instances in which the controller $\hat{K}$ is stable. These measures allow us to precisely quantify the performance of the algorithms relative to the optimal solution. To address the non-stabilizing data-collection policy, we set ${K}_{c}=\alpha I_{3}$, with $\alpha\sim\mathcal{U}\left(-0.1,0\right)$, take $\Sigma_{\eta}=I_{m}$ and fix the reset bound to $d=1000$. Finally, to examine the effect of the rescaling step, we evaluate R-LSVI both with and without the rescaling factor $\alpha_{t}$.
Figure 1. Convergence and stability of the nominal control (PI and VI), O-LSPI, LSPI, R-LSVI (with and without discount) and policy gradient algorithms on the data center cooling benchmark problem as the sample size increases. (Left) Cost relative error. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
Figure 1 illustrates the convergence behavior as a function of the sample size. The left panel displays the convergence of the cost relative error, while the right panel shows the fraction of stable controllers. Our numerical experiments show that all VI algorithms eventually converge to the optimal solution with a cost relative error close to $\approx 10^{-5}$ (policy gradient $\approx 10^{-2}$). However, only VI-based procedures achieve full stabilization at sample size of about $10^{5}$, whereas PI-based procedures do not produce stabilizing controllers. This highlights that VI does not require a stabilizing initial policy. Despite being a PI method, O-LSPI behaves like VI due to its inner VI loop.
We observe that the rescaling step plays a central role for stability when confronted with uneven data collection. Indeed, R-LSVI attains full stability for a sample size of $10^{4}$, whereas its non-rescaling version requires about $10^{5}$. Note that nominal VI converges slightly faster than the ADP methods, consistent with [40, 1, 58, 59].
Finally, the policy gradient baseline benefits from a favorable initialization $\hat{K}_{0}$ and access to exact gradients; therefore, it serves as an optimistic benchmark for model-free methods. Nevertheless, R-LSVI matches or exceeds its performance.
For the remainder of our empirical analysis, we adopt a stabilizing data-collection policy that allows us to focus on other aspects of our procedure. Under this policy, the sampled data are more homogeneous and concentrate around zero, eliminating the need for the reset and rescaling steps.
5.2. Adaptivity under Non-Quadratic Costs
We present an example illustrating the adaptability of the algorithms to changes in the cost function for the benchmark problem. Specifically, we use the same system dynamics (73)–(74), but change the cost function to
$\tilde{c}(x_{t},u_{t})\triangleq x_{t}^{\top}Sx_{t}+(|u_{t}|^{\frac{{\kappa}}{2}})^{\top}R|u_{t}|^{\frac{{\kappa}}{2}}$ (77)
where $|\cdot|^{\frac{{\kappa}}{2}}$ is applied elementwise and ${\kappa}\in[1,3]$. Since this cost function is non-quadratic in $u$ for values other than ${\kappa}=2$, the optimal cost is not expected to adhere to the standard form $\mathrm{Tr}\left(C^{\top}P^{*}C\right)$. Consequently, direct comparisons of relative costs are not meaningful in this setting. Instead, we evaluate the cost obtained using a control input $u_{t}=-\hat{K}x_{t}$, where $\hat{K}$ is computed for each method using a sample size of $T=10^{5}$. The choice of $T$ is guided by prior results (Figure 1), which show that this sample size enables most methods to reach $100\%$ stability. To assess adaptability to cost misspecification independently of the non-stabilizing data-collection policy, we set $K_{c}=\alpha I_{3}$ with $\alpha\sim\mathcal{U}(0.1,0.2)$.
Figure 2. Convergence and stability of the nominal control (VI and PI), LSPI, O-LSPI, R-LSVI and policy gradient algorithms on the data center cooling benchmark problem with non-quadratic cost. (Left) Costs as the exponent $\kappa$} varies. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
In Figure 2, we observe that the ADP algorithms consistently outperform nominal model-based control in both adaptability and stability. In the left panel, for ${\kappa}=2$, the nominal control slightly outperforms the other methods, which is consistent with the exact model specification and the findings in Figure 1. However, beyond this point, R-LSVI systematically outperforms all other algorithms, exhibiting superior adaptability. Notably, the right panel shows that for ${\kappa}>2$, nominal model-based control becomes unstable, with stability dropping sharply below $20\%$. In contrast, non-model-based methods maintain robust performance, with R-LSVI achieving the highest stability, remaining above $80\%$ across all values of $\kappa$}.
Remark 9.
We emphasize that, theoretically, the policy gradient method enjoys a significant advantage because it has direct access to the modified cost function and the ability to compute the exact gradient through backpropagation. Thus, it stands the best chance to identify the implicit optimal linear control compared to the other methods. Nonetheless, we observe that in this example R-LSVI delivers the best controllers in terms of cost.
6. An Application to Dynamic Portfolio Allocation
In this section, we apply the R-LSVI algorithm to a practical financial problem, referred to as dynamic portfolio allocation, using a modeling framework similar to Gârleanu and Pedersen [21]. This model admits a standard stochastic LQR representation [3, 46], with the state vector $x_{t}\in\mathbb{R}^{2N+M}$ that is the concatenation of the portfolio weights in each asset at time t1t-1 ($w_{t-1}\in\mathbb{R}^{N}$), mean-reverting strategy signals ($f_{t}\in\mathbb{R}^{M}$) and asset returns between [t1,t)[t-1,t) ($r_{t}\in\mathbb{R}^{N}$) and the control input $u_{t}\in\mathbb{R}^{N}$ representing the change in portfolio weights at time $t$ ($u_{t}\triangleq w_{t}-w_{t-1}$). In this model we assume the asset returns, $r_{t+1}$, are a linear combination of the mean-reverting signal $f_{t}$ leading to the linear dynamic
$A$ $\triangleq\left[\begin{array}[]{lll}I_{N}&\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ 0&0\\ 0&I_{M}-\Phi&0\\ 0&\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \Pi&0\end{array}\right],B\triangleq\left[\begin{array}[]{l}I_{N}\\ 0\\ 0\end{array}\right],C\triangleq\left[\begin{array}[]{ll}0&0\\ \Omega^{\frac{1}{2}}&0\\ 0&\Sigma^{\frac{1}{2}}\end{array}\right]$ (87)
where $\Pi\in\mathbb{R}^{N\times M}$ is the matrix of factor loadings, $\Phi\in\mathbb{R}^{M\times M}$ is the matrix of the mean-reversion coefficients with $\rho(I_{M}-\Phi)<1$, $\Sigma\in\mathbb{S}^{N}_{++}$, $\Omega\in\mathbb{S}^{M}_{++}$, and $\{\epsilon^{r}_{t}\}$ and $\{\epsilon^{f}_{t}\}$ are the covariance matrix of the white noise processes representing return noise and factors innovations.
The investor aims to select a dynamic trading strategy $(u_{t})_{t>0}$ that maximizes long-term average return adjusted for risk (defined as the portfolio variance) after price impact, leading to the following cost matrices
$S\triangleq\left[\begin{array}[]{ccc}\gamma\Sigma&0&-I_{N}\\ 0&\star_{f}&0\\ -I_{N}&0&\star_{r}\end{array}\right],R\triangleq\Lambda$ (91)
where $\gamma>0$ is the risk aversion coefficient and $\Lambda\in\mathbb{S}^{M}_{++}$ is the matrix that quantifies price impact.
We remark that adding quadratic terms in $f_{t}$ and $r_{t}$ to the cost function does not affect the optimal control since these variables are not controllable. This gives us flexibility in choosing the matrices $\star_{f}\in\mathbb{R}^{M\times M}$ and $\star_{r}\in\mathbb{R}^{N\times N}$ so that the conditions $\lambda_{\mathrm{min}}\left(S\right)>0$ is satisfied. In particular, as we demonstrate in the Appendix D, setting $\star_{f}=I_{M}$ and $\star_{r}=2\gamma^{-1}\Sigma^{-1}$ meets this requirement.
In our model, the factors $\{f_{t}\}$ serves as predictors for forecasting future returns. In real-world trading applications, developing such factors, referred to as “alphas,” is a crucial element of successful dynamic portfolio allocation. While the development of predictors is beyond the scope of this article, we simulate real-world alphas by intentionally introducing look-ahead bias by defining one factor for each asset, given by
$f_{i,t}\triangleq\frac{1}{100}\sum_{k=0}^{99}r_{i,t+1-k}\,.$ (92)
This choice of factors yields linear dynamics with a predictive $R^{2}$ of about $1\%$, which is generally considered a meaningful alpha signal in practice. In the quantitative finance literature, this approach of simulating alphas is known as bootstrap or synthetic alphas [39].
In this article, we set $N=3$ and choose the assets to be the three commodity futures contracts for coffee, cocoa and sugar from January 1, 1996 to January 3, 2024. We estimate the covariance matrices $\Sigma$ and $\Omega$ using returns over the full sample, shrinking the correlations $50\%$ toward zero as in [21]. We take $\gamma=30$, $\Lambda=0.03I_{3}$, $x_{0}=0_{9}$, $K_{c}=0_{3\times 9}$ and $\Sigma_{\eta}=I_{3}$.
Figure 3. Convergence and stability of the nominal VI, O-LSPI and R-LSVI algorithms on the dynamic portfolio allocation problem as the sample size increases. (Left) Cost relative error. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
We conduct two experiments for the dynamic portfolio allocation problem: (a) convergence as a function of sample size, and (b) adaptability to non-quadratic cost misspecification, as in Section 5. We exclude the policy gradient method and compare R-LSVI, O-LSPI, and the model-based nominal control. Overall, the resulting convergence and stability behavior is qualitatively similar to that observed in the data center cooling example. Specifically, the left panel of Figure 3 demonstrates that all methods converge as the sample size increases, with the non-model-based procedure achieving a cost relative error near $10^{-5}$ and with nominal control exhibiting slightly faster convergence. Meanwhile, the right panel shows that all methods attain $100\%$ accuracy when the sample size reaches approximately $10^{4}$.
Regarding cost misspecification, the left panel of Figure 4 shows that the nominal model-based control slightly outperforms the non-model-based methods for ${\kappa}\leq 2$, in contrast to the previous example, where this occurred only at ${\kappa}=2$. In the right panel, however, when ${\kappa}>2$, the nominal approach fails to produce a stable control gain, leading to a dramatic increase in final cost. In contrast, the non-model-based methods maintain robust performance, achieving $100\%$ stability with final costs of comparable magnitude.
Figure 4. Convergence and stability of the nominal VI, O-LSPI and R-LSVI algorithms on the dynamic portfolio allocation problem with non-quadratic cost. (Left) Costs as the exponent $\kappa$} varies. Dashed lines represent the median relative error, with the shaded region covering the $25^{th}$ to $75^{th}$ percentiles, estimated from 100100 trajectories. (Right) Frequency of stabilizing controllers found by the algorithms. Only costs corresponding to stable control gains are plotted in the left panel.
7. Conclusions
In this article, we established the convergence, robustness, and stability of value iteration (VI) for data-driven control of stochastic linear quadratic (LQ) systems in discrete time with completely unknown dynamics and costs. Existing ADP methods for discrete-time continuous state-space systems primarily rely on policy iteration (PI), which requires an initial stabilizing policy. In contrast, we develop the first off-policy, robust, non-model-based VI algorithm, termed R-LSVI, which eliminates the need for a prior stabilizing control policy. By leveraging small-disturbance input-to-state stability (ISS), we prove that VI is globally exponentially stable and robust to external disturbances, ensuring that the generated sequence remains within a small neighborhood of the optimal solution when disturbances are sufficiently small.
To validate our theoretical findings, we evaluated R-LSVI against established non-model-based, model-based, and model-free methods using the “data center cooling” benchmark problem. Our empirical results revealed that R-LSVI consistently outperformed other non-model-based approaches in both stability and convergence speed. In addition, R-LSVI demonstrated strong adaptability to non-quadratic costs, underscoring its potential for broader control applications. Finally, we presented a fully data-driven solution to dynamic portfolio allocation, an important problem in quantitative finance, where our non-model-based approach consistently demonstrated reliable convergence and strong adaptability to non-quadratic costs.
Acknowledgments
This work was supported in part by the NYU Research Catalyst Prize (Zhong-Ping Jiang and Petter N. Kolm) and in part by NSF Grant ECCS-2210320 (Zhong-Ping Jiang). Grégoire G. Macqueron gratefully acknowledges support from the Jeffrey and Diane Rosenbluth Fellowship.
Appendix A Auxiliary Results
Throughout the Appendix, we make use of the following notation
$K_{i}^{j}\triangleq(R+B^{\top}P_{i}^{j}B)^{-1}B^{\top}P_{i}^{j}A,\quad A_{i}^{j}\triangleq A-BK_{i}^{j},\quad R_{i}^{j}\triangleq R+B^{\top}P_{i}^{j}B$ (93)
for all $i\in\{\emptyset,\mathbb{Z}^{+}\}$ and $j\in\{*,1,2\}$.
Lemma 1.
  1. (i)
    For any $A\in\mathbb{R}^{n\times m},B\in\mathbb{R}^{m\times p}$ and $C\in\mathbb{R}^{p\times q}$, we have
    $\mathrm{vec}(ABC)=\left(C^{\top}\otimes A\right)\mathrm{vec}(B)\in\mathbb{R}^{np}\,.$ (94)
  2. (ii)
    For any $v_{1}\in\mathbb{R}^{n}$ and $v_{2}\in\mathbb{R}^{m}$, we have
    $v_{1}\otimes v_{2}=\mathrm{vec}(v_{2}v_{1}^{\top})\in\mathbb{R}^{nm}\,.$ (95)
  3. (iii)
    For any $v_{1}\in\mathbb{R}^{n},v_{2}\in\mathbb{R}^{m}$ and $M\in\mathbb{R}^{n\times m}$, we have
    $v_{1}^{\top}Mv_{2}=\left(v_{2}^{\top}\otimes v_{1}^{\top}\right)\mathrm{vec}(M)\in\mathbb{R}\,.$ (96)
  4. (iv)
    For $v\in\mathbb{R}^{n}$ and $S\in\mathbb{S}^{n}$, we have
    $v^{\top}Sv=\mathrm{vec}(vv^{\top})^{\top}\mathrm{vec}(S)=\tilde{v}^{\top}\mathrm{svec}(S)\,.$ (97)
Proof.
See [45]. ∎
Lemma 2.
For any $X\in\mathbb{R}^{m\times n}$ and $Y\in\mathbb{R}^{n\times m}$ such that both $(I_{m}+XY)$ and $(I_{n}+YX)$ are invertible, we have
  1. (i)
    $(I_{m}+XY)^{-1}X=X(I_{n}+YX)^{-1}\,.$
  2. (ii)
    $(I_{m}+XY)^{-1}=I_{m}-(I_{m}+XY)^{-1}XY\,.$
  3. (iii)
    $I_{m}-X(I_{n}+YX)^{-1}Y=(I_{m}+XY)^{-1}\,.$
Proof.
$(I_{m}+XY)^{-1}X$ $=(I_{m}+XY)^{-1}X(I_{n}+YX)(I_{n}+YX)^{-1}$
$=X(I_{n}+YX)^{-1}\,.$ (98)
$(I_{m}+XY)^{-1}$ $=I_{m}-(I_{m}+XY)^{-1}(I_{m}+XY)+(I_{m}+XY)^{-1}$
$=I_{m}-(I_{m}+XY)^{-1}XY\,.$ (99)
$I_{m}-X(I_{n}+YX)^{-1}Y$ $=I_{m}-(I_{m}+XY)^{-1}XY$
$=(I_{m}+XY)^{-1}\,.$ (100)
Lemma 3.
For any $P\in\mathbb{S}^{n}_{+}$, we denote by $\hat{\mathcal{Q}}\left(P\right)\triangleq\mathcal{Q}\left(P\right)+\Delta\mathcal{Q}\left(P\right)$, where $\Delta\mathcal{Q}\left(P\right)\in\mathbb{S}^{n+m}$ is a disturbance, and define $d\triangleq\min(\lambda_{\mathrm{min}}\left(S\right),\lambda_{\mathrm{min}}\left(R\right))$. The following properties hold
  1. (i)
    $\mathcal{Q}\left(P\right)\succeq\mathcal{Q}\left(0_{n\times n}\right)\succ 0$.
  2. (ii)
    If $\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}<d$ then $\hat{\mathcal{Q}}\left(P\right)\in\mathbb{S}^{n+m}_{++}$.
  3. (iii)
    If $\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}<d$ then $\hat{\mathcal{Q}}\left(P\right)]_{uu}$ is invertible.
Proof.
(i) We have that $\mathcal{Q}\left(0_{n\times n}\right)=\begin{bmatrix}S&0\\ 0&R\end{bmatrix}\succ 0$, as $S\in\mathbb{S}^{n}_{++}$ and $R\in\mathbb{S}^{m}_{++}$. Hence,
$\mathcal{G}\left(P\right)=\mathcal{Q}\left(P\right)-\mathcal{Q}\left(0_{n\times n}\right)=\begin{bmatrix}A^{T}PA&A^{T}PB\\ B^{T}PA&B^{T}PB\end{bmatrix}$. (101)
Then, for any vector $x=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix}\in\mathbb{R}^{n+m}$, we have that
$x^{\top}\mathcal{G}\left(P\right)x$ $=\begin{bmatrix}Ax_{1}\\ Bx_{2}\end{bmatrix}^{\top}P\begin{bmatrix}Ax_{1}\\ Bx_{2}\end{bmatrix}\geq 0$ (102)
as $P\in\mathbb{S}^{n}_{+}$.
(ii) Using (i), we have that for any $P\in\mathbb{S}^{n}_{+}$
$\lambda_{\mathrm{min}}\left(\hat{\mathcal{Q}}(P)\right)$ $\geq\lambda_{\mathrm{min}}\left(\mathcal{Q}\left(P\right)\right)-\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}$ (103)
$\geq\lambda_{\mathrm{min}}\left(\mathcal{Q}\left(0_{n\times n}\right)\right)-d$ (104)
$>0.$ (105)
(iii) Recall that $\hat{\mathcal{Q}}\left(P\right)]_{uu}=R+B^{\top}PB{+[\Delta\mathcal{Q}\left(P\right)]_{uu}}$, hence
$\lambda_{\mathrm{min}}\left([\hat{\mathcal{Q}}\left(P\right)]_{uu}\right)$ $\geq\lambda_{\mathrm{min}}\left(R\right)-{\lVert[\Delta\mathcal{Q}\left(P\right)]_{uu}\rVert_{2}}$ (106)
$\geq\lambda_{\mathrm{min}}\left(R\right)-{\lVert\Delta\mathcal{Q}\left(P\right)\rVert_{2}}$ (107)
$>0$ (108)
as $\lambda_{\mathrm{min}}\left(R\right)>0$. ∎
Lemma 4.
For any $P\in\mathbb{S}^{n}_{+}$, $R\in\mathbb{S}^{m}_{++}$ and $B\in\mathbb{R}^{n\times m}$, we have
$P\succeq P-PB(R+B^{\top}PB)^{-1}B^{\top}P\succeq 0\,.$ (109)
Proof.
The first inequality follows directly from $PB(R+B^{\top}PB)^{-1}B^{\top}P\in\mathbb{S}^{n}_{+}$. For the second inequality, we denote by $M=P^{\frac{1}{2}}B$ and $N=MR^{-\frac{1}{2}}$. Then,
$M(R+M^{\top}M)^{-1}M^{\top}$ $=N(I_{m}+N^{\top}N)^{-1}N^{\top}\preceq I_{n}\,.$ (110)
By multiplying inequality (110) by $P^{\frac{1}{2}}$ from both left and right, the proof is complete. ∎
Lemma 5.
Suppose that $P^{i},P^{j}\in\mathbb{S}^{n}_{+}$. Then,
$K^{i}-K^{j}=R^{-1}B^{\top}\Lambda_{i}^{-\top}(P^{i}-P^{j})A^{j}$ (111)
where $\Lambda_{i}=I_{n}+BR^{-1}B^{\top}P^{i}$.
Proof.
By the definition of $K^{j}$,
$K^{j}$ $=R^{-1}(I_{m}+B^{\top}P^{j}BR^{-1})^{-1}B^{\top}P^{j}A=R^{-1}B^{\top}P^{j}\Lambda_{j}^{-1}A.$ (112)
Hence, we have
$K^{i}-K^{j}$ $=R^{-1}B^{\top}(P^{i}-P^{j})\Lambda_{j}^{-1}A$ (113)
$\quad+R^{-1}B^{\top}P^{i}\Lambda_{i}^{-1}A-R^{-1}B^{\top}P^{i}\Lambda_{j}^{-1}A\,.$
Let us focus on the second term $R^{-1}B^{\top}P^{i}\Lambda_{i}^{-1}A-R^{-1}B^{\top}P^{i}\Lambda_{j}^{-1}A\triangleq(*)$
$(*)$ $=R^{-1}B^{\top}P^{i}(\Lambda_{i}^{-1}\Lambda_{j}\Lambda_{j}^{-1}-\Lambda_{i}^{-1}\Lambda_{i}\Lambda_{j}^{-1})A$ (114)
$=R^{-1}B^{\top}P^{i}\Lambda_{i}^{-1}(\Lambda_{j}-\Lambda_{i})\Lambda_{j}^{-1}A$ (115)
$=-R^{-1}B^{\top}P^{i}\Lambda_{i}^{-1}BR^{-1}B^{\top}(P^{i}-P^{j})\Lambda_{j}^{-1}A\,.$ (116)
Finally,
$K^{i}-K^{j}$ $=R^{-1}B^{\top}[I_{n}-P^{i}\Lambda_{i}^{-1}BR^{-1}B^{\top}](P^{i}-P^{j})\Lambda_{j}^{-1}A$
$=R^{-1}B^{\top}\Lambda_{i}^{-\top}(P^{i}-P^{j})\Lambda_{j}^{-1}A$ (117)
where the last equality follows from Lemma 2 (iii). We conclude by observing that $A^{j}={\Lambda_{j}}^{-1}A$ from Lemma 2 (iii). ∎
Lemma 6.
Let $\{P_{i}^{j}\}_{i\in\mathbb{Z}_{+}}$ with $j\in\{1,2\}$ be the sequences of the following Riccati difference equations
$P_{i+1}^{j}$ $=\mathcal{R}(P_{i}^{j},S_{i}^{j}),\quad P_{0}^{j}\in\mathbb{S}^{n}_{+}.$ (118)
Then the sequence $\Delta{P}_{i}=P_{i}^{1}-P_{i}^{2}$ follows the following dynamic
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}[\Delta{P}_{i}-\Delta{P}_{i}B(R+B^{\top}{P}^{1}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A_{i}^{2}+\Delta S_{i}$ (119)
with $\Delta S_{i}=S^{1}_{i}-S^{2}_{i}$.
Proof.
Equation (118) with $j=1$ can be rewritten as
$P_{i+1}^{1}$ $=A^{\top}P_{i}^{1}A+S^{1}_{i}-(K_{i}^{1})^{\top}R_{i}^{1}K_{i}^{1}$ (120)
$=(A_{i}^{2})^{\top}P_{i}^{1}A_{i}^{2}+S^{1}_{i}-(K_{i}^{1})^{T}R_{i}^{1}K_{i}^{1}+(K_{i}^{2})^{\top}B^{\top}P_{i}^{1}A+A^{\top}P_{i}^{1}BK_{i}^{2}$ (121)
$\qquad-(K_{i}^{2})^{\top}(B^{\top}P_{i}^{1}B)K_{i}^{2}$
$=(A_{i}^{2})^{\top}P_{i}^{1}A_{i}^{2}+S^{1}_{i}-(K_{i}^{1})^{T}R_{i}^{1}K_{i}^{1}+(K_{i}^{2})^{\top}R_{i}^{1}K_{i}^{1}+(K_{i}^{1})^{\top}R_{i}^{1}K_{i}^{2}$ (122)
$\qquad-(K_{i}^{2})^{\top}(R_{i}^{1}-R)K_{i}^{2}\,.$
In addition, we can rewrite (118) with $j=2$ as
$P_{i+1}^{2}$ $=(A_{i}^{2})^{\top}P_{i}^{2}A_{i}^{2}+S^{2}_{i}+(K_{i}^{2})^{\top}RK_{i}^{2}\,.$ (123)
Subtracting (123) from (120) yields
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}-(K_{i}^{1}-K_{i}^{2})^{\top}(R+B^{\top}P_{i}^{1}B)(K_{i}^{1}-K_{i}^{2}).$ (124)
Finally, by applying Lemma 5 to (124) we obtain
$\Delta{P}_{i+1}$ $=(A_{i}^{2})^{\top}\Delta{P}_{i}A_{i}^{2}+\Delta S_{i}-(A_{i}^{2})^{\top}\Delta{P}_{i}B(R+B^{\top}{P}^{1}_{i}B)^{-1}B^{\top}\Delta{P}_{i}A_{i}^{2}.$ (125)
Lemma 7 (Matrix Differentiation, [45]).
Let $M\in\mathbb{R}^{n\times n}$ and $S\subset\mathbb{R}^{n\times q}$ be an open subset such that for every $P\in S$, $(I_{n}+PM)$ is invertible. Define the matrix-valued function $U:S\rightarrow\mathbb{R}^{n\times n}$ by $U(P)\triangleq(I_{n}+PM)^{-1}$. Then $U$ is differentiable on $S$ and the Jacobian is the $n^{2}\times n^{2}$ matrix
$\frac{d\mathrm{vec}(U(P))}{d\mathrm{vec}(P)}=(I_{n}+MP)^{-\top}\otimes(I_{n}+PM)^{-1}\,.$ (126)
Proof.
We note that for any matrix function $F:\mathbb{R}^{n\times p}\rightarrow\mathbb{R}^{m\times q}$, $d\mathrm{vec}(F(X))=\mathrm{vec}(dF(X))$. Then
$d\left((I_{n}+PM)^{-1}P\right)$ $=d\left((I_{n}+PM)^{-1}\right)P+(I_{n}+PM)^{-1}d\left(P\right)$
$=(I_{n}+PM)^{-1}d\left(P\right)(I_{n}+PM)^{-1}+(I_{n}+PM)^{-1}d\left(P\right)$
$=(I_{n}+PM)^{-1}d\left(P\right)\left[I_{n}-M(I_{n}+PM)^{-1}P\right]$
$=(I_{n}+PM)^{-1}d\left(P\right)(I_{n}+MP)^{-1}$ (127)
where the last two equalities follow from Lemma 2 (iii). By applying $\mathrm{vec}(\cdot)$ on both sides of the identity above and using Lemma 1 (i) the proof is complete. ∎
Appendix B Proof of Proposition 1
Controllability: Consider the non-resetting chain
$X_{t+1}=LX_{t}+D\xi_{t+1},\quad\xi_{t}\underset{iid}{\sim}\mathcal{N}(0_{m+p},I_{m+p})$ (128)
with $L=A-BK_{c}$ and $D=(B\Sigma_{\eta}^{1/2}\penalty 10000\ C)$. Without loss of generality, we assume that $(L,D)$ is controllable. Since $x_{0}=0_{n}$, each uncontrollable state component is identically equal to zero for all $t$. Their invariant distribution is therefore the point mass $\delta_{0}$.
Irreducibility: Define the $t$-step reachability Gramian
$W_{t}=\sum_{s=0}^{t-1}L^{s}DD^{\top}(L^{s})^{\top}.$ (129)
Since $(L,D)$ is controllable, there exists $T\leq n$ such that $\mathrm{rank}(W_{T})=n$ and hence $W_{T}\succ 0$. Let $\bar{\mathcal{B}}_{1}(0_{n})$ be the closed unit ball and $\mathcal{V}\subset\bar{\mathcal{B}}^{\infty}_{d}(0_{n})$ any non-empty open set, define
$E(\mathcal{V})\triangleq\{x_{0}=0_{n},\ x_{T}\in\mathcal{V},\ x_{i}\in\bar{\mathcal{B}}^{\infty}_{d}(0_{n})\ \forall i<T\}.$ (130)
Because the transition density of $(x_{t})$ is Gaussian and $W_{T}\succ 0$, then $\mathbb{P}(E(\mathcal{V}))>0$. Since $(x_{t})_{t>0}$ and $(X_{t})_{t>0}$ coincide while inside $\bar{\mathcal{B}}^{\infty}_{d}(0_{n})$, the resetting chain $(x_{t})_{t>0}$ can reach any non-empty open set of $\bar{\mathcal{B}}^{\infty}_{d}(0_{n})$ in a finite number of steps $T<n$ from $0_{n}$ with strictly positive probability.
Positive Recurrence: Let $p\triangleq\mathbb{P}(\|D\xi_{t}\|_{2}>2d)$. Since $DD^{\top}\neq 0$, we have $p>0$, independent of $t$ and the current state $x_{t}$. Let $\tau=\inf\{t\geq 0:\|x_{t}\|_{\infty}>d\}$. At each step, the probability of exiting the ball (and thus resetting to 0) is at least $p$, hence
$\mathbb{P}(\tau>k)\leq(1-p)^{k}\text{ and }\mathbb{E}[\tau]\leq\frac{1}{p}<\infty.$ (131)
Thus the reset mechanism defines a regenerative Markov chain with finite mean cycle length and atom $\{0_{n}\}$, implying positive recurrence.
Hence, the regenerative Markov chain admits a unique invariant probability measure $\pi$
$\pi(A)\;=\;\frac{\mathbb{E}\left[\sum_{t=0}^{\tau}\mathbbm{1}_{\{x_{t}\in A\}}\right]}{\mathbb{E}[\tau]},\quad A\subset\bar{\mathcal{B}}^{\infty}_{d}(0_{n})\,.$ (132)
Appendix C Exact VI
Lemma 8.
If $\lVert\Delta\rVert_{\infty}<\lambda_{\mathrm{min}}\left(S\right)$, then $\hat{P}_{i}\succeq 0$ for all $i\in\mathbb{Z}_{+}$.
Proof.
The proof is by induction. Note that $\lVert\Delta\rVert_{\infty}=\sup\{\lVert\Delta_{i}\rVert,i\in\mathbb{Z}_{+}\}<\lambda_{\mathrm{min}}\left(S\right)$ implies that $S+\Delta_{i}\succ 0$ for all $i\in\mathbb{Z}_{+}$. Suppose $\hat{P}_{i}\succeq 0$ for some $i\in\mathbb{Z}_{+}$. Since
$\hat{P}_{i+1}$ $=\mathcal{R}(\hat{P}_{i},S+\Delta_{i})=A^{\top}\left(\hat{P}_{i}-\hat{P}_{i}B(R+B^{\top}\hat{P}_{i}B)^{-1}B^{\top}\hat{P}_{i}\right)A+S+\Delta_{i}$ (133)
it follows that $\hat{P}_{i+1}\succeq 0$ by Lemma 4. As $\hat{P}_{0}\succeq 0$, the proof is complete. ∎
Lemma 9.
Consider the exact value iteration
${P}_{i+1}$ $=\mathcal{R}\left({P}_{i},S\right),\quad{P}_{0}\in\mathbb{S}^{n}_{+}.$ (134)
If $\Delta P_{0}\triangleq P_{0}-P^{*}\in\mathcal{P}\triangleq\{P\in\mathbb{S}^{n}:\lVert P\rVert_{{2}}\leq\frac{1-\lVert A^{*}\rVert_{{2}}}{\lVert BR^{-1}B^{\top}\rVert_{{2}}\lVert A^{*}\rVert_{{2}}}\}$, then for all $i\in\mathbb{Z}_{+}$
$\lVert\Delta P_{i}\rVert_{{2}}\leq\lVert A^{*}\rVert_{{2}}^{i}\lVert\Delta P_{0}\rVert_{{2}}\,.$ (135)
Proof.
By observing that $\Delta P_{i}=P_{i}-P^{*}$ and applying Lemma 6 and Lemma 2 (iii), we have
$\Delta{P}_{i+1}$ $=(A^{*})^{\top}\Delta{P}_{i}[I_{n}-B(R+B^{\top}{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A^{*}$ (136)
$=(A^{*})^{\top}\Delta{P}_{i}[I_{n}-BR^{-1}B^{\top}(I_{n}+{P}_{i}BR^{-1}B^{\top})^{-1}\Delta{P}_{i}]A^{*}.$ (137)
Taking the norm of (137) yields
$\lVert\Delta{P}_{i+1}\rVert_{{2}}$ $\leq\lVert A^{*}\rVert_{{2}}\lVert\Delta{P}_{i}\rVert_{{2}}(1+\lVert BR^{-1}B^{\top}\rVert_{{2}}\lVert\Delta{P}_{i}\rVert_{2})\lVert A^{*}\rVert_{{2}}.$ (138)
If $\Delta P_{i}\in\mathcal{P}$, we have $\lVert\Delta{P}_{i+1}\rVert_{{2}}\leq\lVert A^{*}\rVert_{{2}}\lVert\Delta{P}_{i}\rVert_{{2}}$, and the proof is complete. ∎
C.1. Proof of Proposition 2
The proof is by induction. Clearly, $P_{i}^{1}\succeq P_{i}^{2}$ for $i=0$. Assume $P_{i}^{1}\succeq P_{i}^{2}$ for some $i\in\mathbb{Z}_{+}$. Using Lemma 6 and that $R+B^{\top}P^{1}_{i}B=R^{2}_{i}+B^{\top}\Delta{P}_{i}B$, we obtain
$\Delta{P}_{i+1}$ $=\Delta S_{i}+(A_{i}^{2})^{\top}[\Delta{P}_{i}-\Delta{P}_{i}B(R^{2}_{i}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A_{i}^{2}.$ (139)
By applying Lemma 4, we have that $\Delta{P}_{i+1}\succeq 0$, i.e. $P_{i+1}^{1}\succeq P_{i+1}^{2}$.
C.2. Proof of Proposition 3
The proof follows from the implicit function theorem. We observe that
$\mathcal{R}\left(P,S\right)$ $=A^{\top}[I_{n}-PBR^{-1}(I_{m}+B^{\top}PBR)^{-1}B^{\top}]PA+S$
$=A^{\top}(I_{n}+PBR^{-1}B^{\top})^{-1}PA+S$ (140)
where the last equality follows from Lemma 2. Noting that we can rewrite equation (49) as $0=\mathcal{R}\left(P,S+\delta I_{n}-P\right)$, define
$f(\mathrm{vec}(P);\delta)$ $\triangleq\mathrm{vec}(\mathcal{R}\left(P,S+\delta I_{n}-P\right))$ (141)
$=A^{\top}\otimes A^{\top}\mathrm{vec}((I_{n}+PBR^{-1}B^{\top})^{-1}P)$ (142)
$\qquad+\mathrm{vec}(S+\delta I_{n})-\mathrm{vec}(P)\,.$
Let $M=BR^{-1}B^{\top}$. Using Lemma 7 and that $(I_{n}+MP)^{-\top}=(I_{n}+PM)^{-1}$ for $P$, $M\in\mathbb{S}^{n}$, we have
$\frac{d\mathrm{vec}((I_{n}+PM)^{-1}P)}{d\mathrm{vec}(P)}=(I_{n}+PM)^{-1}\otimes(I_{n}+PM)^{-1}\,.$ (143)
Hence,
$\frac{df(\mathrm{vec}(P);\delta)}{d\mathrm{vec}(P)}=A(P)^{\top}\otimes A(P)^{\top}-I_{n^{2}}$ (144)
where $A(P)\triangleq(I_{n}+MP)^{-1}A$. We apply Lemma 2 (iii) to rewrite $A(P)$ as
$A(P)$ $=A-BK(P)$ (145)
where $K(P)=(R+B^{T}PB)^{-1}B^{\top}PA$. Therefore, for $P=P^{*}(\delta)$ with $\delta\in(-\lambda_{\mathrm{min}}\left(S\right),\infty)$, $A\left(P^{*}(\delta)\right)$ is Schur stable (see, [44, Section 2.4] for details). Since $\lVert A\left(P^{*}(\delta)\right)\rVert<1$, the left hand side of (144) is invertible. Consequently, by the implicit function theorem $P^{*}(\delta)$ is continuously differentiable with respect to $\delta$.
C.3. Proof of Theorem 2
Let $\Delta P_{i}=P_{i}-P^{*}$. Then, Using Lemma 6 and that $R+B^{\top}P_{i}B=R^{*}+B^{\top}\Delta{P}_{i}B$, we obtain
$\Delta{P}_{i+1}$ $=(A^{*})^{\top}[\Delta{P}_{i}-\Delta{P}_{i}B(R^{*}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]A^{*}\,.$ (146)
The statement in the theorem is obviously true if $\lVert A^{*}\rVert_{2}=0$. Next, we assume $\lVert A^{*}\rVert_{2}\neq 0$ and consider the statement in three cases: ① $\Delta{P}_{0}$ is positive semi-definite, i.e. $P_{0}\succeq P^{*}\succeq 0$; ② $\Delta{P}_{0}$ is negative semi-definite, i.e. $0\preceq P_{0}\preceq P^{*}$; and ③ $\Delta{P}_{0}$ is indefinite, i.e. $\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)>0$ and $\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)<0$.
$\Delta{P}_{0}\succeq 0$: First, we show by induction that if $\Delta{P}_{0}\succeq 0$ then $\Delta{P}_{i}\succeq 0$ for all $i\in\mathbb{Z}_{+}$. Indeed, assume $\Delta{P}_{i}\succeq 0$ for some $i\in\mathbb{Z}_{+}$. Since $\Delta{P}_{i}\succeq[\Delta{P}_{i}-\Delta{P}_{i}B(R^{*}+B^{\top}\Delta{P}_{i}B)^{-1}B^{\top}\Delta{P}_{i}]\succeq 0$ by Lemma 4, we have that $\Delta{P}_{i+1}\succeq 0$. Taking the norm of both sides of (146) results in
$\lVert\Delta{P}_{i+1}\rVert_{2}\leq\lVert A^{*}\rVert_{2}^{2}\lVert\Delta{P}_{i}\rVert_{2}.$ (147)
Consequently, we obtain
$\lVert\Delta{P}_{i}\rVert_{2}\leq(\lVert A^{*}\rVert_{2}^{2})^{i}\lVert\Delta{P}_{0}\rVert_{2}\leq\beta(\lVert\Delta{P}_{0}\rVert_{2},i).$ (148)
$\Delta{P}_{0}\preceq 0$: Obviously, $P^{*}\preceq\Delta{P}_{0}\preceq 0$ then the Theorem holds for $\Delta P_{0}\in\mathcal{P}\cap\{P\in\mathbb{S}^{n}|-P^{*}\preceq P\preceq 0\}$ by Lemma 9. Next, we consider $\Delta P_{0}\in\mathcal{P}^{c}\cap\{P\in\mathbb{S}^{n}|-P^{*}\preceq P\preceq 0\}$, where $\mathcal{P}^{c}$ is the complement of $\mathcal{P}$ over $\mathbb{S}^{n}$. Let $\{P_{i}^{0}\}_{i\in\mathbb{Z}_{+}}$ denote the sequence generated by the exact VI with the initial condition $P_{0}^{0}=0$, and $\Delta P_{i}^{0}=P_{i}^{0}-P^{*}$. By [7, Proposition 4.4.1], $\Delta P_{i}^{0}$ is nondecreasing with respect to $i$ and $\lim_{i\to\infty}\Delta P_{i}^{0}=0$. Let $i_{0}$ denote the smallest index $i$ such that $P_{i}^{0}\in\mathcal{P}$. Since $\Delta P_{0}^{0}\preceq\Delta P_{0}\preceq 0$, by Lemma 8, $\Delta P_{i_{0}}^{0}\preceq\Delta P_{i_{0}}\preceq 0$. Hence, $\lVert\Delta P_{i_{0}}\rVert_{2}\leq\lVert\Delta P_{i_{0}}^{0}\rVert_{2}$ and so $\Delta P_{i}\in\mathcal{P}$ for all $i\geq i_{0}$. For $i\geq i_{0}$, by Lemma 9, we have
$\lVert\Delta P_{i}\rVert_{2}\leq\lVert A^{*}\rVert^{i-i_{0}}_{2}\lVert\Delta P_{i_{0}}\rVert_{2}.$ (149)
For any $0\leq i\leq i_{0}$, we have
$\lVert\Delta P_{i}\rVert_{2}\leq\lVert P^{*}\rVert_{2}$ $\leq\frac{\lVert P^{*}\rVert_{2}\lVert BR^{-1}B^{\top}\rVert_{2}\lVert A^{*}\rVert_{2}}{1-\lVert A^{*}\rVert_{2}}\lVert\Delta P_{0}\rVert_{2}$
$\leq\frac{\lVert P^{*}\rVert_{2}\lVert BR^{-1}B^{\top}\rVert_{2}\lVert A^{*}\rVert_{2}}{(1-\lVert A^{*}\rVert_{2})\lVert A^{*}\rVert^{i_{0}}_{2}}\lVert A^{*}\rVert_{2}^{i}\lVert\Delta P_{0}\rVert_{2}$
$=\alpha\lVert A^{*}\rVert_{2}^{i}\lVert\Delta P_{0}\rVert_{2}.$ (150)
It follows from (149) and (C.3) that
$\lVert\Delta P_{i}\rVert_{2}\leq{\alpha}\lVert A^{*}\rVert_{2}^{i}\lVert\Delta P_{0}\rVert_{2}=\beta(\lVert\Delta P_{0}\rVert_{2},i),\quad\forall i\in\mathbb{Z}_{+}.$ (151)
$\Delta{P}_{0}$ is indefinite: Obviously, $\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)I_{n}\preceq\Delta{P}_{0}\preceq\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)I_{n}$. Let $\{P^{+}_{i}\}_{i\in\mathbb{Z}_{+}}$ and $\{P^{-}_{i}\}_{i\in\mathbb{Z}_{+}}$ be the sequences generated by the exact VI with $P^{+}_{0}=\lambda_{\mathrm{max}}\left(\Delta{P}_{0}\right)I_{n}+P^{*}$ and $P^{-}_{0}=\lambda_{\mathrm{min}}\left(\Delta{P}_{0}\right)I_{n}+P^{*}$, respectively. Since $P_{0}^{+}-P^{*}\succeq 0$, it follows from Case ① that $\lVert P_{i}^{+}-P^{*}\rVert_{2}\leq\beta(\lambda_{\mathrm{max}}\left(P_{0}-P^{*}\right),i)$. In addition, since $P^{-}_{0}-P^{*}\preceq 0$, by Case ②, we have that $\lVert P_{i}^{-}-P^{*}\rVert_{2}\leq$ $\beta(-\lambda_{\mathrm{min}}\left(P_{0}-P^{*}\right),i)$. Since $P_{0}^{-}\preceq P_{0}\preceq P_{0}^{+}$, it follows from Proposition 2,
$P_{i}^{-}-P^{*}\preceq P_{i}-P^{*}\preceq P_{i}^{+}-P^{*}$ (152)
for all $i\in\mathbb{Z}_{+}$. Consequently, $\lambda_{\mathrm{min}}\left(P_{i}^{-}-P^{*}\right)\leq\lambda_{\mathrm{min}}\left(P_{i}-P^{*}\right)\leq 0\leq\lambda_{\mathrm{max}}\left(P_{i}-P^{*}\right)\leq\lambda_{\mathrm{max}}\left(P_{i}^{+}-P^{*}\right)$. Hence, we have
$\lVert P_{i}-P^{*}\rVert_{2}$ $=\max\{|\lambda_{\mathrm{min}}\left(P_{i}-P^{*}\right)|,|\lambda_{\mathrm{max}}\left(P_{i}-P^{*}\right)|\}$
$\leq\max\{\beta(-\lambda_{\mathrm{min}}\left(P_{0}-P^{*}\right),i),\beta(\lambda_{\mathrm{max}}\left(P_{0}-P^{*}\right),i)\}$
$=\beta(\lVert P_{0}-P^{*}\rVert_{2},i)$ (153)
and the proof is complete.
Appendix D Dynamic Portfolio Allocation
We demonstrate that the dynamic portfolio allocation problem introduced in Section 6 is a well-defined stochastic LQR problem.
Lemma 10.
For $N,M\in\mathbb{Z}_{+}$, $\gamma>0$ and $\Sigma\in\mathbb{S}^{N}_{++}$, we have
$G=\left[\begin{array}[]{ccc}\gamma\Sigma&0&-I_{N}\\ 0&I_{M}&0\\ -I_{N}&0&2\gamma^{-1}\Sigma^{-1}\end{array}\right]\in\mathbb{S}^{2N+M}_{++}\,.$ (154)
Proof.
By construction $G\in\mathbb{S}^{2N+M}$, so it remains to show that the spectrum of $G$ satisfies $\sigma(G)\subset\mathbb{R}_{+}^{*}$. Defining $\tilde{\Sigma}\triangleq\gamma\Sigma$, we note that $\tilde{\Sigma},\tilde{\Sigma}^{-1}\in\mathbb{S}^{N}_{++}$ are simultaneously diagonalizable. As the eigenvalues are invariant by change of basis, without loss of generality we consider the matrix
$G=\left[\begin{array}[]{ccc}D&0&-I_{N}\\ 0&I_{M}&0\\ -I_{N}&0&2D^{-1}\end{array}\right]\in\mathbb{S}^{2N+M}$ (155)
where $D=\operatorname{diag}(d_{1},\dots,d_{N})$ with $d_{i}=\gamma\sigma_{i}^{2}>0$ and $\{\sigma_{i}^{2}\}_{i=1}^{N}$ are the strictly positive eigenvalues of $\Sigma$.
To determine the eigenvalues of $G$, we need to solve the system
$\left[\begin{array}[]{ccc}D&0&-I_{N}\\ 0&I_{M}&0\\ -I_{N}&0&2D^{-1}\end{array}\right]\left[\begin{array}[]{c}x\\ y\\ z\end{array}\right]=\lambda\left[\begin{array}[]{c}x\\ y\\ z\end{array}\right]$ (156)
where $x,z\in\mathbb{R}^{N},y\in\mathbb{R}^{M}$ and $x^{\top},y^{\top},z^{\top}]^{\top}\neq 0_{2N+M}$. This is equivalent to solving the system
$\begin{cases}d_{i}x_{i}-z_{i}&=\lambda{x_{i}},\quad\forall i\in\{1,\ldots,N\}\\ y_{j}&=\lambda y_{j},\quad\forall j\in\{1,\ldots,M\}\\ 2d_{i}^{-1}z_{i}-x_{i}&=\lambda z_{i},\quad\forall i\in\{1,\ldots,N\}\,.\end{cases}$ (157)
Since we only require the eigenvalues $\lambda$, it suffices to consider the following cases:
$\lambda=0$: Then $0_{2N+M}$ is the only solution. Hence $0\notin\sigma(G)$.
$\lambda=1$: Then $x^{\top},y^{\top},z^{\top}]^{\top}=[0_{N}^{\top},y^{\top},0_{N}^{\top}]^{\top}$ where $y\neq 0_{M}$. Hence $1\in\sigma(G)$.
$\lambda\neq 1$: Then $y=0_{M}$ and the system reduces to
$\begin{cases}z_{i}&=(d_{i}-\lambda)x_{i},\quad\forall i\in\{1,\ldots,N\}\\ x_{i}&=(2d_{i}^{-1}-\lambda)(d_{i}-\lambda)x_{i},\quad\forall i\in\{1,\ldots,N\}\,.\end{cases}$ (158)
We are looking for $x\neq 0_{N}$ (otherwise this leads to $x^{\top},y^{\top},z^{\top}]^{\top}=0_{2N+M}$). Let us denote $i_{0}\in\{1,\cdots,N\}$ an entry such that $x_{i_{0}}\neq 0$. Then $\lambda$ is the positive solution of
$\lambda^{2}-(2d_{i}^{-1}+d_{i})\lambda+1=0$ (159)
given by $\lambda=\frac{1}{2}\Big((2d_{i}^{-1}+d_{i})\pm\sqrt{(2d_{i}^{-1})^{2}+(d_{i})^{2}}\Big)>0$, where the strict positivity follows from $(a+b)^{2}>a^{2}+b^{2}$ for $a,b>0$. Hence $\sigma(G)\subset\mathbb{R}_{+}^{*}$. ∎
Proposition 6.
Define $A,B,S$ as in (87)–(91) with $\star_{f}=I_{M}$ and $\star_{r}=2\gamma^{-1}\Sigma^{-1}$, then
  1. (i)
    $S\in\mathbb{S}^{N}_{++}\,.$
  2. (ii)
    $(A,B)$ is stabilizable.
Proof.
  1. (i)
    This is a direct application of Lemma 10.
  2. (ii)
    We need to prove that $\exists K\in\mathbb{R}^{N\times{(2N+M)}}$ s.t. $\rho\left(A-BK\right)<1$. Setting $K=[I_{N},0_{M\times M},0_{N\times N}]$, we obtain
    $A-BK=\left(\begin{array}[]{ccc}0&0&0\\ 0&I_{M}-\Phi&0\\ 0&\Pi&0\end{array}\right)$ (163)
    and hence $\sigma(A-BK)=\{0\}\bigcup\sigma(I_{M}-\Phi)$, where by assumption $\rho(I_{M}-\Phi)<1$.